### v1

# Recurrence Relations and Dispersive Techniques for Precision Multi-Loop Calculations

Abstract
Ab initio predictions of two-loop electroweak contributions to observables are increasingly essential for precision collider experiments, yet their evaluation remains very challenging. We connect recurrence techniques and dispersive method in order to evaluate complex multi-loop Feynman diagrams. By expressing multi-point Passarino-Veltman functions in a two-point basis and using shifted space-time dimensions with recurrence relations, we minimize the number of required dispersive integrals. This approach reduces computation time and enables a precise and efficient analysis of one- and two-loop diagrams.
I Introduction
Modern collider and low-energy precision programs are driving sub-percent uncertainties, demanding ab initio, two-loop electroweak (EW) predictions for multi-scale, multi-leg processes. To meet this precision frontier, theoretical methods must evolve in parallel with experiments. Flagship measurements include MOLLER (MOLLER, ) at Jefferson Lab (low-$Q^{2}$ determination of $\sin{\theta}_{W}^{2}$ from parity-violating Møller scattering), P2 (P2, ) at MESA (proton weak charge), and Belle II at KEK SuperKEKB (precision flavour and CP-violation studies). Upcoming programs such as SoLID-PVDIS (SOLID, ) at JLab and the Electron-Ion Collider (EIC) (EIC, ) at BNL will further require comprehensive higher-order theory predictions across multiple scattering channels. The accurate theoretical description of electroweak processes has long been a cornerstone of precision tests of the Standard Model (SM). Over the past four decades, the field has progressed from the first analytical one-loop formalisms to modern semi-analytical and numerical two-loop frameworks capable of supporting sub-percent experimental precision. This evolution reflects both major theoretical advances in multi-loop quantum-field-theory techniques and the rising experimental demands of collider and low-energy programs. The systematic treatment of loop corrections in the electroweak theory was established in the late 1970s and 1980s, culminating in the Passarino-Veltman (PV) tensor-reduction method. The PV algorithm provided a general prescription for expressing one-loop tensor integrals in terms of scalar functions, establishing the algebraic foundation for all subsequent higher-order calculations (PV1979, ).
During the following decade, Refs. (Hollik1990, ; Denner1993, ) produced comprehensive reviews and explicit one-loop calculations relevant to LEP precision physics, codifying gauge-invariant renormalization schemes and parameter definitions. These works became standard references for both theoretical and computational approaches to radiative corrections. The extension to two-loop order demanded a deeper algebraic understanding of Feynman integrals. Refs. (Tarasov96, ; Tarasov97, ) introduced dimension-recurrence and propagator-power-reduction identities that relate integrals in $d$ and $d\pm 2$ dimensions. These relations enable systematic reduction of higher-rank and higher-dimensional integrals to a minimal set of master integrals. This formalism remains foundational for modern two-loop reduction algorithms and underpins most symbolic-manipulation packages used today. Concurrently, in (PLB91, ) a complementary algebraic framework was developed for reducing tensor Feynman integrals to scalar ones using dimension shifting recurrence relations.
Analytic evaluation of increasingly complex diagrams soon became infeasible due to the proliferation of mass scales, external invariants, and threshold singularities. This challenge prompted the development of semi-analytical and numerical approaches such as sector-decomposition methods, differential-equation (DE) systems for master integrals, and dispersion-relation techniques (Lee, ; SmrBook, ; Kotikov, ). A major step was the differential-equation approach for master integrals, in which systems of linear equations in kinematic invariants are solved either analytically in canonical ($\varepsilon$-form) basis or numerically using series expansions. In (Wasser, ; Henn, ) this framework was refined by introducing canonical-basis and uniform-weight formulations, improving both analytic transparency and numerical stability.
In (AA-1, ; AA-2, ; AA-3, ; AA-4, ; AA-5, ), it was developed a comprehensive two-loop framework, specifically for polarized Møller scattering: an essential channel for future parity-violation experiments. The group systematically advanced from reducible two-loop and quadratic one-loop contributions to two-loop irreducible self-energies, vertex and box calculations. These results quantified higher-order electroweak effects in parity-violating asymmetries, establishing reliable theoretical uncertainties at the sub-percent level-critical for the forthcoming MOLLER experiment.
Beyond differential-equation and sector-decomposition strategies, a distinct line of development has been the dispersive approach developed in (AA1, )-(AB19a, ). This methodology expresses multi-point Passarino-Veltman functions in a two-point-function basis, replacing sub-loop insertions with effective propagators represented by dispersion integrals. The result is a semi-analytical bridge between purely analytic amplitude reductions and purely numerical integration techniques. The dispersive framework preserves key analytic properties — threshold behavior, unitarity cuts, and gauge invariance — while reducing computational demands. It is especially attractive for low-energy observables where delicate cancellations between diagrams require high numerical precision. Furthermore, real-experiment implementation often involves acceptance and energy-threshold cuts that significantly affect radiative corrections; the dispersive formulation allows these to be incorporated naturally at the numerical-integration stage.
Parallel developments in the phenomenological sector culminated in partial two-loop electroweak predictions for key observables. In Refs. (Dubovyk1, ) and (Freitas1a, ) produced the full set of fermionic and bosonic two-loop corrections to Z-boson observables, while (Erler, ) addressed hadronic effects in Møller scattering at NNLO. More recently, (Schw, ) achieved the analytic evaluation of electroweak double-box integrals relevant for Møller processes.
These advances collectively establish a robust infrastructure for high-precision electroweak phenomenology. A comprehensive review on updated measurements and higher-order theoretical corrections is available at (Erler1, ) and (PDG2024, ). Despite impressive progress, complete NNLO electroweak results remain available only for selected processes due to the technical difficulty of two-loop calculations. Closed-fermion-loop NNLO corrections have been achieved for several key observables, but the general problem of fully automated two-loop amplitude generation, reduction, and evaluation remains open. Semi-numerical strategies, such as the dispersive and differential-equation methods, have proven especially effective, balancing analytical control with numerical tractability. The growing complexity of precision calculations arises from the need to handle diagrams with both more external legs and additional loop orders. Each loop introduces new mass and momentum scales, overlapping ultraviolet and infrared divergences, and complicated threshold structures, dramatically increasing algebraic and numerical challenges. At one loop, PV-style analytic reductions allow tensor integrals to be expressed through a small set of scalar functions, often in closed form. At two loops and beyond, however, the explosion of topologies and kinematic invariants typically renders complete analytic evaluation impractical. To address these challenges, the dispersive method reformulates a multi-loop integral as a sequence of nested one-loop integrals via spectral representations. This effectively transforms a two-loop problem into integrals over well-behaved spectral densities, isolating singular behavior in analytically controlled functions that can be integrated numerically with high stability. By combining dimensional shifting, recurrence relations, and dispersion theory, the approach achieves algebraic reduction to a minimal set of independent integrals while maintaining analytic transparency.
The present study extends our previous dispersive framework by incorporating shifted-dimension tensor decomposition and dimension-lowering recurrence relations directly into the dispersive representation. These relations, originally formulated within Tarasov dimension-recursion approach, are adapted here to operate on spectral integrals, leading to a minimal set of independent dispersive building blocks. This algebraic reduction significantly decreases computational time while maintaining high numerical precision. We demonstrate the applicability of this formalism to one-loop self-energy, triangle, and box diagrams, as well as two-loop example in which one-loop sub-block is represented through corresponding dispersive integral. The results confirm that the dispersive-recurrence combination provides a stable and robust framework for the precision electroweak calculations. This methodology therefore represents a major step toward an automated, high-precision, ab initio framework for two-loop quantum-field-theory calculations-one capable of supporting the next generation of parity-violation and flavour-physics experiments, including MOLLER, P2, Belle II, and the future programs at EIC.
II Methodology
II.1 Tensor decomposition
The tensor $N$-point function of an arbitrary rank $M$ is defined as
$T^{(N)}_{\mu_{1}\ldots\mu_{M}}$ $\!\!=\!\!$ $\frac{\mu^{4-D}\;e^{\gamma_{E}(4-D)/2}}{{\rm i}\pi^{D/2}}\int{\rm d}^{D}q\;$ (1)
$\times\frac{q_{\mu_{1}}\ldots q_{\mu_{M}}}{\left[q^{2}-m_{1}^{2}\right]\left[(k_{1}\!+\!q)^{2}-m_{2}^{2}\right]\left[(k_{1}\!+\!k_{2}\!+\!q)^{2}-m_{3}^{2}\right]\ldots\left[(k_{1}\!+\ldots+\!k_{N-1}\!+\!q)^{2}-m_{N}^{2}\right]},\hskip 14.22636pt$
where $D=4-2\varepsilon$ is the space-time dimension ($\varepsilon$ is the standard dimensional regularization parameter) and $\mu$ is the mass scale. $T^{(N)}_{\mu_{1}\ldots\mu_{M}}$ can be decomposed in terms of the scalar Passarino-Veltman functions $Z_{0\ldots 01\ldots 12\ldots 2\ldots(N-1)\ldots(N-1)}$ as (see, e.g., in Ref. LoopTools )
$T^{(N)}_{\mu_{1}\ldots\mu_{M}}$ $=$ $\mathop{\sum_{l,n_{1},n_{2}}}_{2l+n_{1}+n_{2}+\ldots+n_{N-1}=M}\left\{[g]^{l}[k_{1}]^{n_{1}}[k_{1}+k_{2}]^{n_{2}}\ldots[k_{1}+k_{2}+\ldots+k_{N-1}]^{n_{N-1}}\right\}_{\mu_{1}\ldots\mu_{M}}\;$ (2)
$\hskip 99.58464pt\times Z_{\underset{2l}{\underbrace{\scriptstyle{0...0}}}\;\underset{n_{1}}{\underbrace{\scriptstyle{1...1}}}\;\underset{n_{2}}{\underbrace{\scriptstyle{2...2}}}\;\ldots\underset{n_{N-1}}{\underbrace{\scriptstyle{(N-1)...(N-1)}}}\;}\;,$
where $\left\{[g]^{l}[k_{1}]^{n_{1}}[k_{1}+k_{2}]^{n_{2}}\ldots[k_{1}+k_{2}+\ldots+k_{N-1}]^{n_{N-1}}\right\}_{\mu_{1}\ldots\mu_{M}}$ is the symmetrized tensor structure containing $l$ metric tensors $g$, $n_{1}$ vectors $k_{1}$, $n_{2}$ vectors $k_{1}+k_{2}$, $\ldots$, and $n_{N-1}$ vectors $k_{1}+\ldots+k_{N-1}$ ($2l+n_{1}+n_{2}+\ldots+n_{N-1}=M$).
In the notation of Ref. PLB91 , let us define the scalar $N$-point integral as
$J^{(N)}(D;\nu_{1},\nu_{2},\ldots,\nu_{N})=\int\frac{{\rm d}^{D}q}{\left[(p_{1}+q)^{2}-m_{1}^{2}\right]^{\nu_{1}}\left[(p_{2}+q)^{2}-m_{2}^{2}\right]^{\nu_{2}}\ldots\left[(p_{N}+q)^{2}-m_{N}^{2}\right]^{\nu_{N}}}\;$ (3)
(it is clear that this scalar integral depends only on the squared momenta $(p_{i}-p_{j})^{2}$ with $i<j<N$). Similarly, the $N$-point tensor integral of an arbitrary rank $M$ is defined as
$J^{(N)}_{\mu_{1}\ldots\mu_{M}}(D;\nu_{1},\nu_{2},\ldots,\nu_{N})=\int{\rm d}^{D}q\;\frac{q_{\mu_{1}}\ldots q_{\mu_{M}}}{\left[(p_{1}+q)^{2}-m_{1}^{2}\right]^{\nu_{1}}\left[(p_{2}+q)^{2}-m_{2}^{2}\right]^{\nu_{2}}\ldots\left[(p_{N}+q)^{2}-m_{N}^{2}\right]^{\nu_{3}}}\;.$ (4)
Then the general tensor decomposition formula (see Eq. (11) of PLB91 ) yields
$J^{(N)}_{\mu_{1}\ldots\mu_{M}}(D;\nu_{1},\nu_{2},\ldots\nu_{N})$ $\!\!\!\!=\!\!\!\!$ $\mathop{\sum_{\lambda,\kappa_{1},\ldots,\kappa_{N}}}_{2\lambda+\kappa_{1}+\ldots+\kappa_{N}=M}\!\!\!\!\!\left(-\frac{1}{2}\right)^{\lambda}\;\left(\prod\limits_{i=1}^{N}(\nu_{i})_{\kappa_{i}}\right)\left\{[g]^{\lambda}[p_{1}]^{\kappa_{1}}[p_{2}]^{\kappa_{2}}\ldots[p_{N}]^{\kappa_{N}}\right\}_{\mu_{1}\ldots\mu_{M}}\;\hskip-14.22636pt$ (5)
$\times\pi^{\lambda-M}\;J^{(N)}\left(D+2(M-\lambda);\nu_{1}+\kappa_{1},\nu_{2}+\kappa_{2},\ldots,\nu_{N}+\kappa_{N}\right)\;,\hskip 14.22636pt$
where $(\nu)_{\kappa}\equiv\Gamma(\nu+\kappa)/\Gamma(\nu)$ is the Pochhammer symbol, and the scalar integrals $J^{(N)}$ occurring on the r.h.s. have shifted space-time dimension value $D+2(M-\lambda)$.
Figure 1: The one-loop NN-point diagram in the notation corresponding to tensors $T^{(N)}_{\mu_{1}\ldots\mu_{M}}$
Figure 2: The one-loop NN-point diagram in the notation corresponding to tensors $J^{(N)}_{\mu_{1}\ldots\mu_{M}}$
Comparing the definitions of $T^{(N)}_{\mu_{1}\ldots\mu_{M}}$ (given in Eq. (1)) and $J^{(N)}_{\mu_{1}\ldots\mu_{M}}$ (given in Eq. (4)) we get the following connection formula:
$T^{(N)}_{\mu_{1}\ldots\mu_{M}}=\frac{\mu^{4-D}e^{\gamma_{E}(4-D)/2}}{{\rm i}\pi^{D/2}}\;J^{(N)}_{\mu_{1}\ldots\mu_{M}}(D;1,\ldots,1)\Bigr|_{p_{1}=0,\;p_{2}=k_{1},\ldots,p_{N}=k_{1}+\ldots+k_{N-1}}\;.$ (6)
Considering Eq. (5) in the case $p_{1}=0,\;p_{2}=k_{1},\ldots,p_{N}=k_{1}+\ldots+k_{N-1}$ we see that only the term with $\kappa_{1}=0$ contributes,
$J^{(N)}_{\mu_{1}\ldots\mu_{M}}(D;\nu_{1},\nu_{2},\ldots\nu_{N})\Bigr|_{p_{1}=0,\;p_{2}=k_{1},\ldots,p_{N}=k_{1}+\ldots+k_{N-1}}$
$\hskip 14.22636pt=\mathop{\sum_{\lambda,\kappa_{2},\ldots,\kappa_{N}}}_{2\lambda+\kappa_{2}+\ldots+\kappa_{N}=M}\!\!\!\!\left(-\frac{1}{2}\right)^{\lambda}\!\left(\prod\limits_{i=2}^{N}(\nu_{i})_{\kappa_{i}}\!\right)\left\{[g]^{\lambda}[k_{1}]^{\kappa_{2}}[k_{1}\!+\!k_{2}]^{\kappa_{3}}\ldots[k_{1}\!+\ldots+\!k_{N-1}]^{\kappa_{N}}\right\}_{\mu_{1}\ldots\mu_{M}}\;$
$\hskip 28.45274pt\times\pi^{\lambda-M}\;J^{(N)}\left(D+2(M\!-\!\lambda);\nu_{1},\nu_{2}+\kappa_{2},\ldots,\nu_{N}+\kappa_{N}\right)\Bigr|_{p_{1}=0,\;p_{2}=k_{1},\ldots,p_{N}=k_{1}+\ldots+k_{N-1}}.\hskip 14.22636pt$ (7)
To compare with $T^{(N)}_{\mu_{1}\ldots\mu_{M}}$ we need to put $\nu_{1}=\nu_{2}=\ldots=\nu_{N}=1$, $\lambda=l$, $\kappa_{2}=n_{1}$, $\kappa_{3}=n_{2}$, $\ldots$, $\kappa_{N}=n_{N-1}$,
$\hskip-14.22636ptJ^{(N)}_{\mu_{1}\ldots\mu_{M}}(D;1,\ldots,1)\Bigr|_{p_{1}=0,\;p_{2}=k_{1},\ldots,p_{N}=k_{1}+\ldots+k_{N-1}}$
$=\mathop{\sum_{l,n_{1},\ldots,n_{N-1}}}_{2l+n_{1}+\ldots+n_{N-1}=M}\left(-\frac{1}{2}\right)^{l}\;\left(\prod\limits_{i=1}^{N-1}{n_{i}}!\right)\left\{[g]^{l}[k_{1}]^{n_{1}}[k_{1}\!+\!k_{2}]^{n_{2}}\ldots[k_{1}\!+\ldots+\!k_{N-1}]^{n_{N-1}}\right\}_{\mu_{1}\ldots\mu_{M}}\;$
$\hskip 14.22636pt\times\pi^{-l-n_{1}-\ldots-n_{N-1}}J^{(N)}\left(D\!+\!2l\!+\!2n_{1}+\ldots+\!2n_{N-1};1,1\!+\!n_{1},\ldots,1\!+\!n_{N-1}\right)\Bigr|_{p_{1}=0,\;p_{2}=k_{1},\ldots,p_{N}=k_{1}+\ldots+k_{N-1}}.\hskip-56.9055pt$
In this way, we arrive at
$Z_{\underset{2l}{\underbrace{\scriptstyle{0...0}}}\;\underset{n_{1}}{\underbrace{\scriptstyle{1...1}}}\;\underset{n_{2}}{\underbrace{\scriptstyle{2...2}}}\;\ldots\underset{n_{N-1}}{\underbrace{\scriptstyle{(N-1)...(N-1)}}}}$
$\hskip 14.22636pt=\frac{\mu^{4-D}e^{\gamma_{E}(4-D)/2}}{{\rm i}\pi^{l+n_{1}+\ldots+n_{N-1}+D/2}}\left(\prod\limits_{i=1}^{N-1}{n_{i}}!\right)\left(-\frac{1}{2}\right)^{l}$
$\hskip 28.45274pt\times J^{(N)}\left(D\!+\!2l\!+\!2n_{1}+\ldots+\!2n_{N-1};1,1\!+\!n_{1},\ldots,1\!+\!n_{N-1}\right)\Bigr|_{p_{1}=0,\;p_{2}=k_{1},\ldots,p_{N}=k_{1}+\ldots+k_{N-1}}.\hskip 14.22636pt$ (9)
To be consistent with the standard Passarino-Veltman notations, for $N=2,3,4,5,\ldots$ the notation $Z$ should be replaced by $B$, $C$, $D$, $E$, etc.
For the two-point case ($N=2$) we get
$B_{\stackrel{{\scriptstyle\underbrace{\scriptstyle{0\ldots 0}}}}{{2l}}\;\stackrel{{\scriptstyle\underbrace{\scriptstyle{1\ldots 1}}}}{{n}}}=\frac{\mu^{4-D}e^{\gamma_{E}(4-D)/2}n!}{{\rm i}\pi^{l+n+D/2}}\left(-\frac{1}{2}\right)^{l}J^{(2)}\left(D+2l+2n;1,1+n\right)\Bigr|_{p_{1}=0,\;p_{2}=k_{1}}\;.$ (10)
If we have only one external momentum we will usually suppress its index, $k_{1}=k$.
For the three-point case ($N=3$) we get
$C_{\underset{2l}{\underbrace{\scriptstyle{0...0}}}\;\underset{n_{1}}{\underbrace{\scriptstyle{1...1}}}\;\underset{n_{2}}{\underbrace{\scriptstyle{2...2}}}}$ $=$ $\frac{\mu^{4-D}e^{\gamma_{E}(4-D)/2}n_{1}!n_{2}!}{{\rm i}\pi^{l+n_{1}+n_{2}+D/2}}\left(-\frac{1}{2}\right)^{l}$ (11)
$\times J^{(3)}\left(D+2l+2n_{1}+2n_{2};1,1+n_{1},1+n_{2}\right)\Bigr|_{p_{1}=0,\;p_{2}=k_{1},\;p_{3}=k_{1}+k_{2}}\;,$
where the scalar integral $J^{(3)}$ on the r.h.s. depends on the following momentum invariants: $(p_{1}-p_{2})^{2}=k_{1}^{2}$ (for the incoming momentum opposite to the line with $m_{3}$), $(p_{2}-p_{3})^{2}=k_{2}^{2}$ (for the incoming momentum opposite to the line with $m_{1}$) and $(p_{3}-p_{1})^{2}=(k_{1}+k_{2})^{2}$ (for the incoming momentum opposite to the line with $m_{2}$).
In the occurring three-point integral we can combine any pair of denominators by using the Feynman parametrization trick, e.g.,
$\hskip-14.22636ptJ^{(3)}\left(D+2l+2n_{1}+2n_{2};1,1+n_{1},1+n_{2}\right)\Bigr|_{p_{1}=0,\;p_{2}=k_{1},\;p_{3}=k_{1}+k_{2}}$
$=\int\frac{{\rm d}^{D+2l+2n_{1}+2n_{2}}q}{\left[q^{2}-m_{1}^{2}\right]\left[(k_{1}+q)^{2}-m_{2}^{2}\right]^{1+n_{1}}\left[(k_{1}+k_{2}+q)^{2}-m_{3}^{2}\right]^{1+n_{2}}}$
$=\frac{(n_{1}+n_{2}+1)!}{n_{1}!n_{2}!}\int\limits_{0}^{1}{\rm d}x\;x^{n_{1}}\bar{x}^{n_{2}}\int\frac{{\rm d}^{D+2l+2n_{1}+2n_{2}}q}{\left[q^{2}-m_{1}^{2}\right]\left[(k_{1}+\bar{x}k_{2}+q)^{2}-xm_{2}^{2}-\bar{x}m_{3}^{2}+x\bar{x}k_{2}^{2}\right]^{2+n_{1}+n_{2}}}$
$=\frac{(n_{1}\!+\!n_{2}\!+\!1)!}{n_{1}!n_{2}!}\int\limits_{0}^{1}{\rm d}x\;x^{n_{1}}\bar{x}^{n_{2}}J^{(2)}(D+2l+2n_{1}+2n_{2};1,2+n_{1}+n_{2})\Bigr|_{\footnotesize{\begin{array}[]{l}p_{1}=0,\;p_{2}=k_{1}+\bar{x}k_{2},\\[-8.53581pt] m_{2}^{2}\leftrightarrow xm_{2}^{2}\!+\!\bar{x}m_{3}^{2}\!-\!x\bar{x}k_{2}^{2}\end{array}}}\;,$ (14)
where $\bar{x}=1-x$.
For the four-point case ($N=4$) we get
$D_{\underset{2l}{\underbrace{\scriptstyle{0...0}}}\;\underset{n_{1}}{\underbrace{\scriptstyle{1...1}}}\;\underset{n_{2}}{\underbrace{\scriptstyle{2...2}}}\;\underset{n_{3}}{\underbrace{\scriptstyle{3...3}}}}$ $=$ $\frac{\mu^{4-D}e^{\gamma_{E}(4-D)/2}n_{1}!n_{2}!n_{3}!}{{\rm i}\pi^{l+n_{1}+n_{2}+n_{3}+D/2}}\left(-\frac{1}{2}\right)^{l}$ (15)
$\hskip-113.81102pt\times J^{(4)}\left(D\!+\!2l\!+\!2n_{1}\!+\!2n_{2}\!+\!2n_{3};1,1\!+\!n_{1},1\!+\!n_{2},1\!+\!n_{3}\right)\Bigr|_{p_{1}=0,\;p_{2}=k_{1},\;p_{3}=k_{1}+k_{2},\;p_{4}=k_{1}+k_{2}+k_{3}}\;.\hskip 14.22636pt$
In the occurring four-point integral we can combine any triple of denominators by using the Feynman parametrization trick, e.g.,
$\hskip-14.22636ptJ^{(4)}\left(D+2l+2n_{1}+2n_{2}+2n_{3};1,1+n_{1},1+n_{2},1+n_{3}\right)\Bigr|_{p_{1}=0,\;p_{2}=k_{1},\;p_{3}=k_{1}+k_{2},\;p_{4}=k_{1}+k_{2}+k_{3}}$
$=\!\int\!\frac{{\rm d}^{D+2l+2n_{1}+2n_{2}+2n_{3}}q}{\left[q^{2}-m_{1}^{2}\right]\left[(k_{1}+q)^{2}-m_{2}^{2}\right]^{1+n_{1}}\left[(k_{1}+k_{2}+q)^{2}-m_{3}^{2}\right]^{1+n_{2}}\left[(k_{1}+k_{2}+k_{3}+q)^{2}-m_{4}^{2}\right]^{1+n_{3}}}$
$=\frac{(n_{1}+n_{2}+n_{3}+2)!}{n_{1}!n_{2}!n_{3}!}\int\limits_{0}^{1}\int\limits_{0}^{1}{\rm d}x\;{\rm d}y\;x^{n_{1}}\bar{x}^{n_{2}+n_{3}+1}y^{n_{2}}\bar{y}^{n_{3}}$
$\hskip 8.53581pt\times\!\int\!\frac{{\rm d}^{D+2l+2n_{1}+2n_{2}+2n_{3}}q}{\left[q^{2}\!-\!m_{1}^{2}\right]\left[(k_{1}\!+\!\bar{x}k_{2}\!+\!\bar{x}\bar{y}k_{3}\!+\!q)^{2}\!-\!xm_{2}^{2}\!-\!\bar{x}ym_{3}^{2}\!-\!\bar{x}\bar{y}m_{4}^{2}\!+\!x\bar{x}(k_{2}\!+\!\bar{y}k_{3})^{2}\!+\!\bar{x}y\bar{y}k_{3}^{2}\right]^{3+n_{1}+n_{2}+n_{3}}}$
$=\frac{(n_{1}+n_{2}+n_{3}+2)!}{n_{1}!n_{2}!n_{3}!}\int\limits_{0}^{1}\int\limits_{0}^{1}{\rm d}x\;{\rm d}y\;x^{n_{1}}\bar{x}^{n_{2}+n_{3}+1}y^{n_{2}}\bar{y}^{n_{3}}$
$\hskip 8.53581pt\times J^{(2)}(D\!+\!2l\!+\!2n_{1}\!+\!2n_{2}\!+\!2n_{3};1,3\!+\!n_{1}\!+\!n_{2}\!+\!n_{3})\Bigr|_{\footnotesize{\begin{array}[]{l}p_{1}=0,\;p_{2}=k_{1}+\bar{x}k_{2}+\bar{x}\bar{y}k_{3},\\[-8.53581pt] m_{2}^{2}\leftrightarrow xm_{2}^{2}\!+\!\bar{x}ym_{3}^{2}\!+\!\bar{x}\bar{y}m_{4}^{2}\!-\!x\bar{x}(k_{2}+\bar{y}k_{3})^{2}\!-\!\bar{x}y\bar{y}k_{3}^{2}\end{array}}}\;,$ (18)
where $\bar{x}=1-x$ and $\bar{y}=1-y$.
II.2 Recurrence relations and the momentum expansion
In the two-point case, according to Eq. (10), we need to calculate the integrals $J^{(2)}\left(D+2l+2n;1,1+n\right)$ with $l\geq 0$ and $n\geq 0$. Using Feynman parameters we can express higher functions in terms of the integrals $J^{(2)}$. In the three- and four-point cases, according to Eqs. (II.1) and (II.1), we need to calculate the integrals $J^{(2)}(D+2l+2n_{1}+2n_{2};1,2+n_{1}+n_{2})$ and $J^{(2)}(D+2l+2n_{1}+2n_{2}+2n_{3};1,3+n_{1}+n_{2}+n_{3})$, respectively. In general, for an $N$-point function we would need to deal with the integrals
$J^{(2)}(D+2l+2n;1,N-1+n),\quad{\rm with}\quad n=n_{1}+\ldots n_{N-1}\;.$ (20)
To decrease the index (power of propagator) N1+nN-1+n, we can use the recurrence relation (87) for the case $\nu_{1}=1$,
$J^{(2)}(D+2;1,\nu_{2}+1)$ $=$ $-\frac{\pi}{2\nu_{2}k^{2}}\left[(k^{2}+m_{1}^{2}-m_{2}^{2})J^{(2)}(D;1,\nu_{2})\right$. (21)
$\hskip 45.5244pt\left.-J^{(2)}(D;1,\nu_{2}-1)+J^{(2)}(D;0,\nu_{2})\right].$
Whenever one of the indices on the r.h.s. becomes zero (like, e.g., in $J^{(2)}(D;0,\nu_{2})$), this is a tadpole integral which can be expressed in terms of $J^{(2)}(D;0,1)$ or $J^{(2)}(D;0,1)$ using Eqs. (83)–(84). After that, for the remaining integrals $J^{(2)}\left(D+2l;1,1\right)$ we can use the recurrence relation (82) for the case $\nu_{1}=\nu_{2}=1$,
$J^{(2)}(D+2;1,1)$ $\!\!\!=\!\!\!$ $-\frac{\pi}{2k^{2}(D-1)}\left[\Delta J^{(2)}(D;1,1)\right$. (22)
$\left.+(k^{2}+m_{1}^{2}-m_{2}^{2})J^{(2)}(D;1,0)+(k^{2}-m_{1}^{2}+m_{2}^{2})J^{(2)}(D;0,1)\right]\;,\hskip 14.22636pt$
with
$\Delta\equiv\Delta(m_{1}^{2},m_{2}^{2},k^{2})=-\lambda(m_{1}^{2},m_{2}^{2},k^{2})=4m_{1}^{2}m_{2}^{2}-(k^{2}-m_{1}^{2}-m_{2}^{2})^{2}\;,$ (23)
where $\lambda(m_{1}^{2},m_{2}^{2},k^{2})$ is the standard notation for the Källen function (other representations of $\Delta$ are collected in Eq. (78)).
Another way to deal with the integrals (20) is to use recurrence relations (76)–(77). In this way, we can bring them to the integrals $J^{(2)}(D+2l+2n;1,1)$ (with the same value of the space-time dimension as the original ones) plus tadpoles, and then apply Eq. (22) as many times as needed. However, this option would involve more steps, and it would produce very cumbersome intermediate expressions because of presence of $\Delta$ in the denominators of (76)–(77). We found the way based on Eqs. (21) and (22) to be more efficient.
Let us first consider the two- and three-point cases (we will discuss the higher cases later). In the two-point case ($N=2$), using $n$ times Eq. (21) and $l$ times Eq. (22), we reduce $J^{(2)}\left(D+2l+2n;1,1+n\right)$ to $J^{(2)}(D;1,1)$ plus tadpoles. In the three-point case ($N=3$), using $n+1$ times Eq. (21) and $l$ times Eq. (22), we reduce $J^{(2)}\left(D+2l+2n;1,2+n\right)$ to $J^{(2)}(D-2;1,1)$ plus tadpoles. If we want to use the same basis for the two-point integrals, we need to apply Eq. (22) one more time (shifting $D\to D-2$).
In this way, in the two- and three-point cases, starting from the integrals (20) with $D=4-2\varepsilon$ we bring them to the basis of $(2-2\varepsilon)$-dimensional integrals. Namely, by using recurrence relations with respect to the powers of propagators $\nu_{i}$ and the space-time dimension $D$, we can express all the relevant integrals $J^{(2)}\left(4-2\varepsilon+2l+2n;1,N-1+n\right)$ (with $N=2$ and $N=3$) in terms of the master integral $J^{(2)}(2-2\varepsilon;1,1)$, as well as the tadpoles
$J^{(2)}(2-2\varepsilon;1,0)=-{\rm i}\pi^{1-\varepsilon}\;\Gamma(\varepsilon)\;(m_{1}^{2})^{-\varepsilon}\quad{\rm and}\quad J^{(2)}(2-2\varepsilon;0,1)=-{\rm i}\pi^{1-\varepsilon}\;\Gamma(\varepsilon)\;(m_{2}^{2})^{-\varepsilon}$ (24)
(see Appendix A for more details). Analytical results for the integral $J^{(2)}(2-2\varepsilon;1,1)$ (including the relevant terms of the $\varepsilon$-expansion) are collected in Appendix B. This procedure provides analytical results for all required Passarino-Veltman functions (10).
When using the recurrence relations (21) and (22) we are getting powers of $k^{2}$ in the denominator. In particular, when reducing $J^{(2)}\left(4-2\varepsilon+2l+2n;1,N-1+n\right)$ (with $N=2$ or $N=3$) to $J^{(2)}(2-2\varepsilon;1,1)$ and tadpoles (24) the maximal power is $(k^{2})^{l+n+1}$, i.e.,
$J^{(2)}\left(4\!-\!2\varepsilon\!+\!2l\!+\!2n;1,N\!-\!1\!+\!n\right)$ $=$ $\frac{\pi^{l+n+1}}{(k^{2})^{l+n+1}}\Bigl[R_{N,l,n}^{(1,1)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;1,1\right)$ (25)
$\hskip 48.36967pt+R_{N,l,n}^{(1,0)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;1,0\right)$
$\hskip 48.36967pt+R_{N,l,n}^{(0,1)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;0,1\right)\Bigr]\;,\hskip 22.76219pt$
where $R_{N,n,l}^{(1,1)}$, $R_{N,n,l}^{(1,0)}$ and $R_{N,n,l}^{(0,1)}$ are algebraic coefficients which are polynomial in $k^{2}$.
To make sure that the resulting expression (25) is not singular as $k^{2}\to 0$, let us employ the small momentum expansion of the integral $J^{(2)}(2-2\varepsilon;1,1)$. According to Eq. (124), the terms of the small-$k^{2}$ expansion of $J^{(2)}(2-2\varepsilon;1,1)$ up to $(k^{2})^{j_{0}}$ can be presented as
$J_{[j_{0}]}^{(2)}(2-2\varepsilon;1,1)$ $\!\!=\!\!$ $\sum\limits_{j=0}^{j_{0}}(k^{2})^{j}\;\frac{(1+\varepsilon)_{j}}{(m_{2}^{2}-m_{1}^{2})^{1+2j}}$ (26)
$\times\Biggl\{J^{(2)}(2-2\varepsilon;0,1)\;\sum\limits_{l=0}^{j}\frac{j!}{l!(j-l)!}\;\frac{(m_{2}^{2})^{l}(m_{1}^{2})^{j-l}}{(1-\varepsilon)_{l}(1+\varepsilon)_{j-l}}\;$
$\hskip 14.22636pt-J^{(2)}(2-2\varepsilon;1,0)\;\sum\limits_{l=0}^{j}\frac{j!}{l!(j-l)!}\;\frac{(m_{1}^{2})^{l}(m_{2}^{2})^{j-l}}{(1-\varepsilon)_{l}(1+\varepsilon)_{j-l}}\;\Biggr\}\;,\hskip 22.76219pt$
so that
$J_{[\infty]}^{(2)}(2-2\varepsilon;1,1)=J^{(2)}(2-2\varepsilon;1,1)\;.$ (27)
If we subtract the expansion (26) from $J^{(2)}(2-2\varepsilon;1,1)$, the difference will be of the order $(k^{2})^{j_{0}+1}$, and it can be presented as
$J^{(2)}(2-2\varepsilon;1,1)-J_{[j_{0}]}^{(2)}(2-2\varepsilon;1,1)=(k^{2})^{j_{0}+1}\bar{J}_{j_{0}+1}^{(2)}(2-2\varepsilon;1,1)\;.$ (28)
In our case we need to put $j_{0}=n+l$. Using Eq. (28) we get
$J^{(2)}(2-2\varepsilon;1,1)=J_{[n+l]}^{(2)}(2-2\varepsilon;1,1)+(k^{2})^{n+l+1}\bar{J}_{n+l+1}^{(2)}(2-2\varepsilon;1,1)\;.$ (29)
Combining Eqs. (25), (26) and (29) we get
$J^{(2)}\left(4\!-\!2\varepsilon\!+\!2l\!+\!2n;1,N\!-\!1\!+\!n\right)$ $\!=\!$ $\frac{\pi^{l+n+1}}{(k^{2})^{l+n+1}}\Bigl[R_{N,l,n}^{(1,1)}(m_{1},m_{2},k^{2},\varepsilon)(k^{2})^{n+l+1}\bar{J}_{n+l+1}^{(2)}\left(2-2\varepsilon;1,1\right)\hskip-22.76219pt$ (30)
$\hskip 48.36967pt+\widetilde{R}_{N,l,n}^{(1,0)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;1,0\right)$
$\hskip 48.36967pt+\widetilde{R}_{N,l,n}^{(0,1)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;0,1\right)\Bigr]\;,\hskip 22.76219pt$
where $\widetilde{R}_{N,l,n}^{(1,0)}$ and $\widetilde{R}_{N,l,n,}^{(0,1)}$ include $R_{N,l,n}^{(1,0)}$ and $R_{N,l,n}^{(0,1)}$ plus the polynomial (in $k^{2}$) contributions coming from $R_{N,l,n}^{(1,1)}J_{[n+l]}^{(2)}(2-2\varepsilon;1,1)$ (see Eq. (26)). The absence of singularities in $k^{2}$ means that in $\widetilde{R}_{N,l,n}^{(1,0)}$ and $\widetilde{R}_{N,l,n}^{(0,1)}$ all the powers of $k^{2}$ less than $n+l+1$ should cancel, so that
$\widetilde{R}_{N,l,n}^{(1,0)}(m_{1},m_{2},k^{2},\varepsilon)$ $=$ $(k^{2})^{l+n+1}\bar{R}_{N,l,n}^{(1,0)}(m_{1},m_{2},k^{2},\varepsilon),$
$\widetilde{R}_{N,l,n}^{(0,1)}(m_{1},m_{2},k^{2},\varepsilon)$ $=$ $(k^{2})^{l+n+1}\bar{R}_{N,l,n}^{(0,1)}(m_{1},m_{2},k^{2},\varepsilon),$ (31)
where $\bar{R}_{N,l,n}^{(1,0)}$ and $\bar{R}_{N,l,n}^{(0,1)}$ are also polynomial in $k^{2}$. In this way, we arrive at
$J^{(2)}\left(4-2\varepsilon+2l+2n;1,N-1+n\right)$ $=$ $\pi^{l+n+1}\Bigl[R_{N,l,n}^{(1,1)}(m_{1},m_{2},k^{2},\varepsilon)\bar{J}_{l+n+1}^{(2)}\left(2-2\varepsilon;1,1\right)$ (32)
$\hskip 42.67912pt+\bar{R}_{N,l,n}^{(1,0)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;1,0\right)$
$\hskip 42.67912pt+\bar{R}_{N,l,n}^{(0,1)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;0,1\right)\Bigr]\;.\hskip 22.76219pt$
In particular, this yields the following result for the function (10):
$B_{\stackrel{{\scriptstyle\underbrace{\scriptstyle{0\ldots 0}}}}{{2l}}\;\stackrel{{\scriptstyle\underbrace{\scriptstyle{1\ldots 1}}}}{{n}}}$ $=$ $\frac{\mu^{2\varepsilon}e^{\gamma_{E}\varepsilon}n!}{{\rm i}\pi^{1-\varepsilon}}\left(-\frac{1}{2}\right)^{l}\Bigl[R_{2,l,n}^{(1,1)}(m_{1},m_{2},k^{2},\varepsilon)\bar{J}_{l+n+1}^{(2)}\left(2-2\varepsilon;1,1\right)$ (33)
$\hskip 85.35826pt+\bar{R}_{2,l,n}^{(1,0)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;1,0\right)$
$\hskip 85.35826pt+\bar{R}_{2,l,n}^{(0,1)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;0,1\right)\Bigr]\;.$
Note that the coefficient functions $R_{N,l,n}^{(1,1)}$, $\bar{R}_{N,l,n}^{(1,0)}$ and $\bar{R}_{N,l,n}^{(0,1)}$ do not have poles in $\varepsilon$ because in the recurrence relations (21) and (22) the only $D$-dependent factor in the denominator is (D1)(D-1) which would never produce $\varepsilon$ for even dimensions.
We can split the function (33) into two parts, the first one containing the 1-point (tadpole-like) integrals $J^{(2)}\left(2-2\varepsilon;1,0\right)$ and $J^{(2)}\left(2-2\varepsilon;0,1\right)$, and the second one involving the genuine (subtracted) 2-point integral $\bar{J}_{n+l+1}^{(2)}\left(2-2\varepsilon;1,1\right)$:
$B_{\stackrel{{\scriptstyle\underbrace{\scriptstyle{0\ldots 0}}}}{{2l}}\;\stackrel{{\scriptstyle\underbrace{\scriptstyle{1\ldots 1}}}}{{n}}}\equiv B_{\{2l,n\}}(k^{2},m_{1}^{2},m_{2}^{2})=B_{\{2l,n\}}^{{\rm 1-point}}(k^{2},m_{1}^{2},m_{2}^{2})+B_{\{2l,n\}}^{{\rm 2-point}}(k^{2},m_{1}^{2},m_{2}^{2}),$ (34)
with
$B_{\{2l,n\}}^{{\rm 1-point}}(k^{2},m_{1}^{2},m_{2}^{2})$ $=$ $\frac{\mu^{2\varepsilon}e^{\gamma_{E}\varepsilon}n!}{{\rm i}\pi^{1-\varepsilon}}\left(-\frac{1}{2}\right)^{l}\Bigl[\bar{R}_{2,l,n}^{(1,0)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;1,0\right)$ (35)
$\hskip 93.89409pt+\bar{R}_{2,l,n}^{(0,1)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;0,1\right)\Bigr],\hskip-14.22636pt$
$B_{\{2l,n\}}^{{\rm 2-point}}(k^{2},m_{1}^{2},m_{2}^{2})$ $=$ $\frac{\mu^{2\varepsilon}e^{\gamma_{E}\varepsilon}n!}{{\rm i}\pi^{1-\varepsilon}}\left(-\frac{1}{2}\right)^{l}R_{2,l,n}^{(1,1)}(m_{1},m_{2},k^{2},\varepsilon)\bar{J}_{n+l+1}^{(2)}\left(2-2\varepsilon;1,1\right).$ (36)
Note that all UV-singularities are in $B_{\{2l,n\}}^{{\rm 1-point}}(k^{2},m_{1}^{2},m_{2}^{2})$, namely in the tadpole integrals (24), whereas the term $B_{\{2l,n\}}^{{\rm 2-point}}(k^{2},m_{1}^{2},m_{2}^{2})$ is UV-finite.
For the four-point function the situation is a bit more complicated. Let us start from the integral (20), $J^{(2)}(D+2l+2n;1,3+n)$, and use recurrence relations (21) and (22). If we stop the recurrence procedure when the space-time dimension becomes D2D-2 (i.e., $2-2\varepsilon$) then among the remaining integrals we may have not only $J^{(2)}(2-2\varepsilon;1,1)$, $J^{(2)}(2-2\varepsilon;1,0)$ and $J^{(2)}(2-2\varepsilon;0,1)$, but also $J^{(2)}(2-2\varepsilon;1,2)$ (this happens at $l=0$). The integral $J^{(2)}(2-2\varepsilon;1,2)$ is not independent: using the relation (77) it can be expressed as
$J^{(2)}(D;1,2)$ $=$ $\frac{1}{\Delta}\Bigl[(D-3)(k^{2}+m_{1}^{2}-m_{2}^{2})J^{(2)}(D;1,1)$ (37)
$-(D-2)J^{(2)}(D;1,0)-\frac{(D-2)(k^{2}-m_{1}^{2}-m_{2}^{2})}{2m_{1}^{2}}J^{(2)}(D;0,1)\Bigr]\;.$
If we use Eq. (37) for $J^{(2)}(2-2\varepsilon;1,2)$ we would get for $N=4$ a representation similar to (25), but the occurring coefficient functions $R_{4,l,n}^{(1,1)}$, etc., will not be polynomial in $k^{2}$, because of the presence of $\Delta$ (see Eq. (23)) in their denominators. In this way, we would get rather cumbersome expressions for the higher-order Passarino-Veltman functions.
Another way is to keep the $J^{(2)}(2-2\varepsilon;1,2)$ contributions as an extra term
$R_{N,l,n}^{(1,2)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}(2-2\varepsilon;1,2)$ (38)
in Eq. (25), as well as in Eqs. (30) and (32). For the small-$k^{2}$ expansion we can use the derivative of Eq. (26) w.r.t. $m^{2}$,
$J_{[j_{0}]}^{(2)}(2-2\varepsilon;1,2)=\frac{\partial}{\partial m_{2}^{2}}J_{[j_{0}]}^{(2)}(2-2\varepsilon;1,1)\;,$ (39)
which can be calculated automatically. In this way, for the four-point case we get the following decomposition:
$J^{(2)}\left(4-2\varepsilon+2l+2n;1,3+n\right)$ $=$ $\pi^{l+n+1}\Bigl[R_{4,l,n}^{(1,1)}(m_{1},m_{2},k^{2},\varepsilon)\bar{J}_{l+n+1}^{(2)}\left(2-2\varepsilon;1,1\right)$ (40)
$\hskip 42.67912pt+R_{4,l,n}^{(1,2)}(m_{1},m_{2},k^{2},\varepsilon)\bar{J}_{l+n+1}^{(2)}\left(2-2\varepsilon;1,2\right)$
$\hskip 42.67912pt+\bar{R}_{4,l,n}^{(1,0)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;1,0\right)$
$\hskip 42.67912pt+\bar{R}_{4,l,n}^{(0,1)}(m_{1},m_{2},k^{2},\varepsilon)J^{(2)}\left(2-2\varepsilon;0,1\right)\Bigr]\;.\hskip 22.76219pt$
In the same way, for the five-point function we would get in Eq. (25) an extra term involving $J^{(2)}(2-2\varepsilon;1,3)$, etc.
II.3 The dispersion approach
The subtracted integral $\bar{J}_{n+l+1}^{(2)}\left(2-2\varepsilon;1,1\right)$ can be presented through the dispersive integral as
$\bar{J}_{n+l+1}^{(2)}\left(2-2\varepsilon;1,1\right)=\frac{{\rm i}}{\pi}\;\int\limits_{(m_{1}+m_{2})^{2}}^{\infty}{\rm d}s\;\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}}{s^{n+l+1}\;(s-k^{2}-{\rm i}0)}\;,$ (41)
where (see Eq. (106))
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]_{s}=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\frac{\pi}{\sqrt{-\Delta_{s}}}\left(\frac{s}{-\Delta_{s}}\right)^{\varepsilon}\;$ (42)
(the subscript $s$ means that we substitute $k^{2}\to s$). The first two terms ($\varepsilon^{0}$ and $\varepsilon^{1}$) of the $\varepsilon$-expansion of ${\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]$ are given in Eq. (105). Note that the appearance of the factor $1/s^{n+l}$ in the integrand of Eq. (41) provides better convergence of the dispersive integral. This is another advantage of subtracting the first terms of the Taylor expansion in $k^{2}$.
To derive the dispersive integral representation for $\bar{J}_{n+l+1}^{(2)}\left(2-2\varepsilon;1,2\right)$ we can differentiate Eq. (41) w.r.t. $m_{2}^{2}$,
$\bar{J}_{n+l+1}^{(2)}\left(2-2\varepsilon;1,2\right)=\frac{\partial}{\partial m_{2}^{2}}\bar{J}_{n+l+1}^{(2)}\left(2-2\varepsilon;1,1\right)\;.$ (43)
For the function in the integrand we get
$\frac{\partial}{\partial m_{2}^{2}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}$ $=$ ${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,2)\right]_{s}$ (44)
$=$ $-\frac{(1+2\varepsilon)(s+m_{1}^{2}-m_{2}^{2})}{\Delta_{s}}{\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]_{s}\;.\hskip 22.76219pt$
The same result can be obtained by using Eq. (37) and taking into account that the tadpoles do not contribute to the imaginary part. Note that the limit of integration in (41) also depends on $m_{2}$.
Taking into account that separate terms may have singularities as $s\to(m_{1}+m_{2})^{2}$, let us shift the lower limit by a small positive $\delta$,
$s_{\delta}=(m_{1}+m_{2})^{2}+\delta,$ (45)
and at the end consider the limit $\delta\to 0$. Differentiating Eq. (41) w.r.t. $m_{2}^{2}$ we get
$\bar{J}_{n+l+1}^{(2)}\left(2-2\varepsilon;1,2\right)$ $=$ $\frac{{\rm i}}{\pi}\;\int\limits_{s_{\delta}}^{\infty}\frac{{\rm d}s}{s^{n+l+1}\;(s-k^{2}-{\rm i}0)}\;\frac{\partial}{\partial m_{2}^{2}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}$ (46)
$-\frac{{\rm i}}{\pi}\;\frac{m_{1}+m_{2}}{m_{2}}\;\frac{1}{s_{\delta}^{n+l+1}\;(s_{\delta}-k^{2}-{\rm i}0)}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s_{\delta}}\;.\hskip 22.76219pt$
To proceed, let us use the analytic result given in Eq. (42) to calculate the derivatives
$\frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}\right\}=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\pi\;\left(\frac{\partial\Delta_{s}}{\partial s}\right)\left(\frac{\partial}{\partial\Delta_{s}}(-\Delta_{s})^{-1/2-\varepsilon}\right)\;,$
$\frac{\partial}{\partial m_{2}^{2}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\pi\;s^{\varepsilon}\left(\frac{\partial\Delta_{s}}{\partial m_{2}^{2}}\right)\left(\frac{\partial}{\partial\Delta_{s}}(-\Delta_{s})^{-1/2-\varepsilon}\right)\;.$
Combining these equations we get
$\frac{\partial}{\partial m_{2}^{2}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}=s^{\varepsilon}\frac{\left(\partial\Delta_{s}/\partial m_{2}^{2}\right)}{\left(\partial\Delta_{s}/\partial s\right)}\frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}\right\}\;.$ (47)
Taking into account that
$\frac{\partial\Delta_{s}}{\partial m_{2}^{2}}=2(s+m_{1}^{2}-m_{2}^{2}),\qquad{\rm and}\qquad\frac{\partial\Delta_{s}}{\partial s}=-2(s-m_{1}^{2}-m_{2}^{2})$
we arrive at
$\frac{\partial}{\partial m_{2}^{2}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}=-s^{\varepsilon}\frac{s+m_{1}^{2}-m_{2}^{2}}{s-m_{1}^{2}-m_{2}^{2}}\ \ \frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}\right\}\;.$ (48)
After transforming the derivative w.r.t. $m_{2}^{2}$ into the derivative w.r.t. $s$, we can apply integration by parts to the integral on the r.h.s. of Eq. (46),
$\frac{{\rm i}}{\pi}\;\int\limits_{s_{\delta}}^{\infty}\frac{{\rm d}s}{s^{n+l+1}\;(s-k^{2}-{\rm i}0)}\;\frac{\partial}{\partial m_{2}^{2}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}$
$=-\frac{{\rm i}}{\pi}\;\int\limits_{s_{\delta}}^{\infty}\frac{s^{\varepsilon}\;{\rm d}s}{s^{n+l+1}\;(s-k^{2}-{\rm i}0)}\;\frac{s+m_{1}^{2}-m_{2}^{2}}{s-m_{1}^{2}-m_{2}^{2}}\frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}\right\}$
$=\frac{{\rm i}}{\pi}\;\int\limits_{s_{\delta}}^{\infty}{\rm d}s\;s^{-\varepsilon}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}\frac{\partial}{\partial s}\left[\frac{s^{\varepsilon}}{s^{n+l+1}\;(s-k^{2}-{\rm i}0)}\frac{s+m_{1}^{2}-m_{2}^{2}}{s-m_{1}^{2}-m_{2}^{2}}\right]$
$\hskip 14.22636pt+\frac{{\rm i}}{\pi}\frac{s_{\delta}+m_{1}^{2}-m_{2}^{2}}{s_{\delta}-m_{1}^{2}-m_{2}^{2}}\ \frac{1}{s_{\delta}^{n+l+1}\;(s_{\delta}-k^{2}-{\rm i}0)}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s_{\delta}}\hskip 22.76219pt$ (49)
Recalling that $s_{\delta}=(m_{1}+m_{2})^{2}+\delta$ we can see that in the limit $\delta\to 0$ the last term on the r.h.s. of Eq. (II.3) exactly cancels the non-integral term in Eq. (46). Since the the first (integral) term on the r.h.s. of Eq. (II.3) is finite as $\delta\to 0$, we can put $\delta=0$. In this way we get
$\bar{J}_{n+l+1}^{(2)}\left(2-2\varepsilon;1,2\right)$ $=$ $\frac{{\rm i}}{\pi}\;\int\limits_{(m_{1}+m_{2})^{2}}^{\infty}{\rm d}s\;s^{-\varepsilon}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}$ (50)
$\times\frac{\partial}{\partial s}\left[\frac{s^{\varepsilon}}{s^{n+l+1}\;(s-k^{2}-{\rm i}0)}\ \frac{s+m_{1}^{2}-m_{2}^{2}}{s-m_{1}^{2}-m_{2}^{2}}\right]\;.$
Using Eq. (50) we can also get another representation,
$\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,2\right)$ $=$ $\frac{{\rm i}}{\pi}\!\int\limits_{(m_{1}+m_{2})^{2}}^{\infty}\!\!\!\!\!{\rm d}s\Bigl\{s^{-\varepsilon}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}\!-(k^{2})^{-\varepsilon}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{k^{2}}\Bigr\}$ (51)
$\qquad\qquad\times\frac{\partial}{\partial s}\left[\frac{s^{\varepsilon}}{s^{n+l+1}\;(s-k^{2}-{\rm i}0)}\ \frac{s+m_{1}^{2}-m_{2}^{2}}{s-m_{1}^{2}-m_{2}^{2}}\right]$
$-\frac{{\rm i}}{\pi}\;\frac{m_{1}+m_{2}}{m_{2}}\ \frac{\left[(m_{1}+m_{2})^{2}\right]^{\varepsilon-n-l-1}}{(m_{1}+m_{2})^{2}-k^{2}-{\rm i}0}(k^{2})^{-\varepsilon}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{k^{2}}\;,$
where $k^{2}\leftrightarrow k^{2}+{\rm i}0$, in the same way as in the prescription $1/(s-k^{2}-{\rm i}0)$.
To get an alternative representation, let us start from Eq. (46) and substitute the results (44) for the derivative w.r.t. $m_{2}^{2}$:
$\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,2\right)$ $=$ $-\frac{{\rm i}}{\pi}(1+2\varepsilon)\int\limits_{s_{\delta}}^{\infty}{\rm d}s\;\frac{s+m_{1}^{2}-m_{2}^{2}}{s^{n+l+1}(s-k^{2}-{\rm i}0)\Delta_{s}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}$ (52)
$-\frac{{\rm i}}{\pi}\frac{m_{1}+m_{2}}{m_{2}}\ \frac{1}{s_{\delta}^{n+l+1}(s_{\delta}-k^{2}-{\rm i}0)}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s_{\delta}}\;.\hskip 22.76219pt$
Note that the integral in Eq. (52) is singular as $\delta\to 0$, because
$\Delta_{s}=-\left[s-(m_{1}+m_{2})^{2}\right]\left[s-(m_{1}-m_{2})^{2}\right]=-(s-s_{0})(s-s_{1})\,$, (53)
with $s_{0}\equiv(m_{1}+m_{2})^{2}$, $s_{1}\equiv(m_{1}-m_{2})^{2}$. To separate the finite and the divergent contributions, let us employ the identity
$\frac{1}{s-k^{2}-{\rm i}0}=-\frac{s-s_{\delta}}{(s-k^{2}-{\rm i}0)(s_{\delta}-k^{2}-{\rm i}0)}+\frac{1}{s_{\delta}-k^{2}-{\rm i}0}\,.$ (54)
In this way, we get
$\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,2\right)$ $=$ $\frac{{\rm i}}{\pi}\frac{1+2\varepsilon}{s_{\delta}-k^{2}-{\rm i}0}\int\limits_{s_{\delta}}^{\infty}{\rm d}s\;\frac{(s+m_{1}^{2}-m_{2}^{2})(s-s_{\delta})}{s^{n+l+1}(s-k^{2}-{\rm i}0)\Delta_{s}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}$ (55)
$-\frac{{\rm i}}{\pi}\frac{1+2\varepsilon}{s_{\delta}-k^{2}-{\rm i}0}\int\limits_{s_{\delta}}^{\infty}{\rm d}s\;\frac{s+m_{1}^{2}-m_{2}^{2}}{s^{n+l+1}\Delta_{s}}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}$
$-\frac{{\rm i}}{\pi}\frac{m_{1}+m_{2}}{m_{2}}\ \frac{1}{s_{\delta}^{n+l+1}(s_{\delta}-k^{2}-{\rm i}0)}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s_{\delta}}\;.\hskip 22.76219pt$
The first integral on the r.h.s of Eq. (55) is finite as $\delta\to 0$, so that we can put $\delta=0$ and substitute $(s-s_{\delta})/\Delta_{s}=(s-s_{0})/\Delta_{s}=-1/(s-s_{1})$. To deal with the second integral let us employ the analytic result given in Eq. (42) to calculate the derivative
$\frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}\right\}=\frac{(1+2\varepsilon)(s-m_{1}^{2}-m_{2}^{2})}{\Delta_{s}}\;s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}\,.$
Therefore,
$\frac{1+2\varepsilon}{\Delta_{s}}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}=\frac{s^{\varepsilon}}{s-m_{1}^{2}-m_{2}^{2}}\;\frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}\right\}\,.$
Integrating by parts, we can transform the second integral in Eq. (55) as
$\hskip-28.45274pt-\frac{{\rm i}}{\pi}\frac{1+2\varepsilon}{s_{\delta}-k^{2}-{\rm i}0}\int\limits_{s_{\delta}}^{\infty}{\rm d}s\;\frac{s+m_{1}^{2}-m_{2}^{2}}{s^{n+l+1}\Delta_{s}}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}$ (56)
$=$ $-\frac{{\rm i}}{\pi}\frac{1}{s_{\delta}-k^{2}-{\rm i}0}\int\limits_{s_{\delta}}^{\infty}{\rm d}s\;\frac{s^{\varepsilon}(s+m_{1}^{2}-m_{2}^{2})}{s^{n+l+1}(s-m_{1}^{2}-m_{2}^{2})}\;\frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}\right\}$
$=$ $\frac{{\rm i}}{\pi}\frac{1}{s_{\delta}-k^{2}-{\rm i}0}\ \frac{s_{\delta}+m_{1}^{2}-m_{2}^{2}}{s_{\delta}^{n+l+1}(s_{\delta}-m_{1}^{2}-m_{2}^{2})}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s_{\delta}}$
$+\frac{{\rm i}}{\pi}\frac{1}{s_{\delta}-k^{2}-{\rm i}0}\int\limits_{s_{\delta}}^{\infty}{\rm d}s\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}\ s^{-\varepsilon}\frac{\partial}{\partial s}\left[\frac{s^{\varepsilon}(s+m_{1}^{2}-m_{2}^{2})}{s^{n+l+1}(s-m_{1}^{2}-m_{2}^{2})}\right]\;.\hskip 22.76219pt$
We can see that in the limit $\delta\to 0$ the first (non-integral) term on the r.h.s. of Eq. (56) exactly cancels the non-integral term in Eq. (55). The remaining integrals contibuting to $\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,2\right)$ are finite as $\delta\to 0$. In this way, putting $\delta=0$ we arrive at the following alternative representation:
$\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,2\right)$ $=$ $-\frac{{\rm i}}{\pi}\frac{1+2\varepsilon}{s_{0}\!-\!k^{2}\!-\!{\rm i}0}\int\limits_{s_{0}}^{\infty}{\rm d}s\;\frac{s+m_{1}^{2}-m_{2}^{2}}{s^{n+l+1}(s-s_{1})(s-k^{2}-{\rm i}0)}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}$ (57)
$+\frac{{\rm i}}{\pi}\frac{1}{s_{0}\!-\!k^{2}\!-\!{\rm i}0}\int\limits_{s_{0}}^{\infty}{\rm d}s\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}s^{-\varepsilon}\frac{\partial}{\partial s}\left[\frac{s^{\varepsilon}\;(s+m_{1}^{2}-m_{2}^{2})}{s^{n+l+1}\;(s\!-\!m_{1}^{2}\!-\!m_{2}^{2})}\right].\hskip-28.45274pt$
Using partial fractioning in the denominator of the first integral in Eq. (57) we can get the following representation:
$\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,2\right)$ $=$ $-(1+2\varepsilon)\frac{k^{2}+m_{1}^{2}-m_{2}^{2}}{\Delta}\;\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)$ (58)
$+(1+2\varepsilon)\frac{2m_{1}(m_{1}-m_{2})}{\Delta}\;\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\Bigr|_{k^{2}=s_{1}}$
$+\frac{{\rm i}}{\pi}\frac{1}{s_{0}\!-\!k^{2}\!-\!{\rm i}0}\int\limits_{s_{0}}^{\infty}{\rm d}s\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}s^{-\varepsilon}\frac{\partial}{\partial s}\left[\frac{s^{\varepsilon}\;(s+m_{1}^{2}-m_{2}^{2})}{s^{n+l+1}\;(s\!-\!m_{1}^{2}\!-\!m_{2}^{2})}\right].\hskip-28.45274pt$
Furthermore, using integration by parts we can evaluate the remaining integral in Eq. (58) as
$\hskip-85.35826pt\frac{{\rm i}}{\pi}\int\limits_{s_{0}}^{\infty}{\rm d}s\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}s^{-\varepsilon}\frac{\partial}{\partial s}\left[\frac{s^{\varepsilon}\;(s+m_{1}^{2}-m_{2}^{2})}{s^{n+l+1}\;(s\!-\!m_{1}^{2}\!-\!m_{2}^{2})}\right]$
$=\frac{{\rm i}}{\pi}\frac{\partial}{\partial m_{2}^{2}}\int\limits_{s_{0}}^{\infty}\frac{{\rm d}s}{s^{n+l+1}}\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\right]_{s}$
$=\frac{\partial}{\partial m_{2}^{2}}\bar{J}_{n+l}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\Bigr|_{k^{2}=0}\;,$ (59)
where $\bar{J}_{n+l}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\Bigr|_{k^{2}=0}$ is the $(n+l)$-th coefficient of the small-$k^{2}$ expansion of Eq. (26), namely
$\bar{J}_{n+l}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\Bigr|_{k^{2}=0}$ $\!\!=\!\!$ $\frac{(1+\varepsilon)_{n+l}}{(m_{2}^{2}-m_{1}^{2})^{1+2n+2l}}$ (60)
$\times\Biggl\{J^{(2)}(2-2\varepsilon;0,1)\;\sum\limits_{r=0}^{n+l}\frac{(n+l)!}{r!(n+l-r)!}\;\frac{(m_{2}^{2})^{r}(m_{1}^{2})^{n+l-r}}{(1-\varepsilon)_{r}(1+\varepsilon)_{n+l-r}}\;$
$\hskip 14.22636pt-J^{(2)}(2-2\varepsilon;1,0)\;\sum\limits_{r=0}^{j}\frac{(n+l)!}{r!(n+l-r)!}\;\frac{(m_{1}^{2})^{r}(m_{2}^{2})^{n+l-r}}{(1-\varepsilon)_{r}(1+\varepsilon)_{n+l-r}}\;\Biggr\}\;.\hskip 22.76219pt$
In this way, we get
$\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,2\right)$ $=$ $-(1+2\varepsilon)\;\frac{k^{2}+m_{1}^{2}-m_{2}^{2}}{\Delta}\;\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)$ (61)
$+(1+2\varepsilon)\;\frac{2m_{1}(m_{1}-m_{2})}{\Delta}\;\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\Bigr|_{k^{2}=s_{1}}$
$+\frac{1}{s_{0}\!-\!k^{2}\!-\!{\rm i}0}\;\frac{\partial}{\partial m_{2}^{2}}\bar{J}_{n+l}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\Bigr|_{k^{2}=0}\;.$
or
$\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,2\right)$ $=$ $\frac{1}{s_{0}\!-\!k^{2}\!-\!{\rm i}0}\Biggl\{-(1+2\varepsilon)\;\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)$ (62)
$-(1\!+\!2\varepsilon)\frac{2m_{1}(m_{1}\!-\!m_{2})}{k^{2}-s_{1}}\;\Bigl[\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)-\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\Bigr|_{k^{2}=s_{1}}\Bigr]$
$+\frac{\partial}{\partial m_{2}^{2}}\bar{J}_{n+l}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\Bigr|_{k^{2}=0}\Biggr\}\;.$
The dispersion integral representation for $\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)$ is given in Eq. (41), and the combination of integrals in the second line of Eq. (62) can be presented as
$\frac{1}{k^{2}\!-\!s_{1}}\;\Bigl[\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)-\bar{J}_{n+l+1}^{(2)}\left(2\!-\!2\varepsilon;1,1\right)\Bigr|_{k^{2}=s_{1}}\Bigr]=\frac{{\rm i}}{\pi}\;\int\limits_{s_{0}}^{\infty}{\rm d}s\;\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}}{s^{n+l+1}(s\!-\!k^{2}\!-\!{\rm i}0)(s\!-\!s_{1}\!-\!{\rm i}0)}\;.$
III Numerical Examples
In this section we provide a numerical comparison of three- and four-point functions calculated using techniques outlined in the last chapter and Collier (COL1, )-(COL4, ) numerical library.
Figure 3: Numerical results for the three-point functions $C_{0}$, $C_{1}$, $C_{2}$, $C_{00}$, $C_{11}$ and $C_{12}$ ($k_{2}^{2}=-1.5\penalty 10000\ {\rm GeV}^{2}$, $(k_{1}+k_{2})^{2}=m_{3}^{2}$, $m_{1}=0.5\penalty 10000\ {\rm GeV}$, $m_{2}=1.0\penalty 10000\ {\rm GeV}$, $m_{3}=1.5\penalty 10000\ {\rm GeV}$). The functions $C_{2l,n_{1},n_{2}}$ are defined in Eq. (3). Crossed dots are results based on this work and solid lines are produced from Collier library.
For numerical integration over the dispersive and Feynman parameters we have used Mathematica, GlobalAdaptive method. As it can be seen from Fig. 3, results are in excellent agreement with Collier. Numerical comparison for four-point functions is given in Fig. 4.
Figure 4: Numerical results for the four-point function $D_{0}$ ($k_{2}^{2}=-1.5\penalty 10000\ {\rm GeV}^{2}$, $k_{3}^{2}=-2.5\penalty 10000\ {\rm GeV}^{2}$, $k_{4}^{2}=m_{4}^{2}\penalty 10000\ {\rm GeV}^{2}$, $(k_{1}\cdot k_{3})=4.0\penalty 10000\ {\rm GeV}^{2}$, $(k_{2}\cdot k_{3})=-1.0\penalty 10000\ {\rm GeV}^{2}$ $m_{1}=1.5\penalty 10000\ {\rm GeV}$, $m_{2}=0.5\penalty 10000\ {\rm GeV}$, $m_{3}=2.0\penalty 10000\ {\rm GeV}$, $m_{4}=2.5\penalty 10000\ {\rm GeV}$). The functions $D_{2l,n_{1},n_{2},n_{3}}$ are defined in Eq. (15). Crossed dots are results based on this work and solid line is produced from Collier library.
As in the case of three-point functions, four-point example shows that we have rather good consistency with Collier. At this point we are ready to apply derived many-point functions in dispersive representation to the evaluation of two-loop diagrams.
IV Roadmap to two-Loop calculations
As we can see from the previous chapters we have successfully represented one-loop (up to multiplicity four) integrals with an arbitrary tensor rank using recurrence and dispersive methods. In addition we where able to reduce higher multiplicity PV functions to two-point result. Finally, we have adopted dispersive technique introduced in (AA1, )-(AB19, ) to subtracted two-point functions. Now we have analytical results for PV functions with polynomial terms in $k^{2}$ and dispersive term carrying propagator like structure $\propto\frac{1}{(s-k^{2}-i0)}$. This particular representation is most valuable for applications in two-loop calculations for any possible particle physics models. First, if we consider $\left(j+2\right)$-point Feynman graph as an insertion (index $\left(j+2\right)$ means that we have $j$ number of external and two internal legs in the insertion) into the two-loop topology (see Fig. 5), then polynomial terms in external or second loop momenta will be a part of the numerator algebra and dispersive contribution will be treated as an additional propagator in the second loop integral.
Figure 5: The two-loop NN-point diagram ($k_{1:j}\equiv k_{1}+\ldots+k_{j}$).
Let us define two-loop integral in the similar way as it was done in Eq. (1):
$U_{\mu_{1}\ldots\mu_{M+L}}^{(N)}$ $=\left(\frac{\mu^{4-D}e^{\gamma_{E}(4-D)/2}}{{\rm i}\pi^{D/2}}\right)^{2}\int\int{\rm d}^{D}q_{1}\;{\rm d}^{D}q_{2}$
$\times\frac{{q_{1}}_{\mu_{1}}\ldots{q_{1}}_{\mu_{M}}\;{q_{2}}_{\mu_{M+1}}\ldots{q_{2}}_{\mu_{M+L}}}{\left[q_{1}^{2}-m_{N+2}^{2}\right]\left[q_{2}^{2}-m_{N+3}^{2}\right]\left[\left(q_{1}+q_{2}\right)^{2}-m_{1}^{2}\right]\left[\left(k_{1}+q_{1}+q_{2}\right)^{2}-m_{2}^{2}\right]}$
$\times\frac{1}{\left[\left(k_{1}+k_{2}+q_{1}+q_{2}\right)^{2}-m_{3}^{2}\right]...\left[\left(k_{1}+...+k_{j-1}+q_{1}+q_{2}\right)^{2}-m_{j}^{2}\right]}$
$\times\frac{1}{\left[\left(k_{1}+...+k_{j-1}+q_{1}\right)^{2}-m_{j+1}^{2}\right]...\left[\left(k_{1}+...+k_{N-1}+q_{1}\right)^{2}-m_{N+1}^{2}\right]}.$ (63)
We start our evaluation with the integration over one of the loop momentum. Next step is to apply tensor decomposition, and reduce one-loop insertion of Eq. (63) to $J^{(j+2)}\left(D+2l+2n_{1}+\ldots+2n_{j+1};1,1+n_{1},\ldots,1+n_{j+1}\right)$. After that we can apply Feynman trick to reduce number of the propagators in the first loop integration to two, which will result in the two-point function with one of the propagators in the power of $\left(1+j+n_{1}+\ldots+n_{j+1}\right)$: $J^{(2)}\left(D+2l+2n_{1}+\ldots+2n_{j+1};1,1+j+n_{1}+\ldots+n_{j+1}\right)$. Using recurrence approach we can reduce number of the dimensions to $D=2-2\varepsilon$ and propagator’s power to one, resulting in the insertion expressed as a subtracted UV finite $\bar{J}_{l+n_{1}+\ldots+n_{j+1}}^{(2)}\left(2-2\varepsilon;1,1\right)$ two-point function and two UV divergent tadpoles, $J^{(2)}\left(2-2\varepsilon;1,0\right)$ and $J^{(2)}\left(2-2\varepsilon;0,1\right)$, multiplied by the polynomials in external and second loop momenta. Subtracted $\bar{J}_{l+n_{1}+\ldots+n_{j+1}}^{(2)}\left(2-2\varepsilon;1,1\right)$ can be expressed dispersively (see Eq. (41)), and in the second loop integral we will receive an additional propagator and terms in the numerator expressed as polynomials in the momenta. Second loop integration is now reduced to one-loop integral where we can apply well tested packages, such as X (HP1, ; HP2, ), FeynCalc FeyC1 ; FeyC2 , FormCalc (FormC, ) and Form (Form, ) to complete two-loop evaluation.
Figure 6: Two-loop scalar example.
In order to demonstrate how the outlined roadmap can be applied to two-loop calculations, we chose to consider well known example originally introduced in (Bohm, ). We start our example (see Fig. 6) with two-loop integral
$U^{(2)}$ $=\left(\frac{\mu^{4-D}e^{\gamma_{E}(4-D)/2}}{{\rm i}\pi^{D/2}}\right)^{2}$
$\times\int\int\frac{{\rm d}^{D}q_{1}\;{\rm d}^{D}q_{2}}{\left[q_{1}^{2}-m_{4}^{2}\right]\left[q_{2}^{2}-m_{5}^{2}\right]\left[\left(q_{1}+q_{2}\right)^{2}-m_{1}^{2}\right]\left[\left(q_{1}+q_{2}-k\right)^{2}-m_{2}^{2}\right]\left[\left(q_{1}-k\right)^{2}-m_{3}^{2}\right]},$ (64)
and first perform integration over the loop momentum $q_{2}$. That results in $J^{(3)}\left(D;1,1,1\right)$:
$J^{(3)}\left(D;1,1,1\right)=$ $\int\frac{{\rm d}^{D}q_{2}}{\left[q_{2}^{2}-m_{5}^{2}\right]\left[\left(q_{1}+q_{2}\right)^{2}-m_{1}^{2}\right]\left[\left(q_{1}+q_{2}-k\right)^{2}-m_{2}^{2}\right]}.$ (65)
Using Eq. (II.1), we get
$J^{(3)}\left(D;1,1,1\right)=$ $\int\limits_{0}^{1}{\rm d}x\;J^{(2)}\left(D;1,2\right)\Bigr|_{\footnotesize{\begin{array}[]{l}p_{1}=0,p_{2}=q_{1}-\bar{x}k\\[-8.53581pt] m_{1}^{2}\leftrightarrow m_{5}^{2},m_{2}^{2}\leftrightarrow xm_{1}^{2}+\bar{x}m_{2}^{2}-x\bar{x}k^{2}\end{array}}}$ (68)
$=$ $\int\limits_{0}^{1}{\rm d}x\int\frac{{\rm d}^{D}q_{2}}{\left[q_{2}^{2}-m_{5}^{2}\right]\left[\left(q_{1}-\bar{x}k+q_{2}\right)^{2}-xm_{1}^{2}-\bar{x}m_{2}^{2}+x\bar{x}k^{2}\right]^{2}},$ (69)
which effectively gives us a reduction of the three-point integral to a two-point one. Applying recursive approach outlined in Chapter II, we can lower second power of the last propagator in Eq.(69) and arrive to the subtracted $\bar{J}^{(2)}\left(2-2\varepsilon;1,1\right)$ plus terms containing $J^{(2)}\left(2-2\varepsilon;1,0\right)$ and $J^{(2)}\left(2-2\varepsilon;0,1\right)$ (see Eq. (32)). For the subtracted $\bar{J}^{(2)}\left(2-2\varepsilon;1,1\right)$, we can apply dispersive representation stemming from Eqs. (41)–(42). Finally, for $D=4-2\varepsilon$ and $\varepsilon\to 0$ Eq. (65) can be written in the following form:
$J^{(3)}\left(4-2\varepsilon;1,1,1\right)$ $=\frac{1}{2}\int\limits_{0}^{1}{\rm d}x\left\{\frac{\rm{i}\pi^{2-\epsilon}}{m_{5}^{2}-m_{12x}^{2}}\ln\frac{m_{12x}^{2}}{m_{5}^{2}}+\left[\left(q_{1}-\bar{x}k\right)^{2}+m_{5}^{2}-m_{12x}^{2}\right]\pi\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)\right\},$ (70)
where the function $\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)$ is defined in Eq. (41):
$\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)=\frac{{\rm i}}{\pi}\;\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}}{s\left[\left(q_{1}-\bar{x}k\right)^{2}-s+{\rm i}0\right]}\;.$
Here, $m_{12x}^{2}=xm_{1}^{2}+\bar{x}m_{2}^{2}-x\bar{x}k^{2}$ and the imaginary part of ${\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)$ at $\varepsilon=0$ has a simple structure:
$\rm{Im}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}=-\frac{2\pi^{2}}{\sqrt{\left(s-m_{12x}^{2}-m_{5}^{2}\right)^{2}-4m_{12x}^{2}m_{5}^{2}}},$
In Eq. (70) we did not retain linear in $\varepsilon$ terms, since the insertion $J^{(3)}\left(D;1,1,1\right)$, and the entire two-loop integral are UV-finite. At this point, we are ready to complete integration over second loop momentum $q_{1}$:
$U^{(2)}=$ $\frac{1}{2}\int\limits_{0}^{1}{\rm d}x\Biggl\{\frac{1}{m_{5}^{2}-m_{12x}^{2}}\ln\frac{m_{12x}^{2}}{m_{5}^{2}}\;B_{0}$
$-\frac{1}{\pi}\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}{\rm d}s\;\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}}{s}\left[\left(s+m_{5}^{2}-m_{12x}^{2}\right)C_{0}+B_{0}\right]\Biggr\},$ (71)
where in the second-loop integration we have used usual PV functions without dispersive representation. In Eq. (71) the three-point function has the following arguments: $C_{0}\equiv C_{0}\left(k^{2},x^{2}k^{2},\bar{x}^{2}k^{2},m_{4}^{2},m_{3}^{2},s\right)$ (here we have used the following mapping of arguments for $C_{0}\equiv C_{0}\left(k_{1}^{2},k_{2}^{2},(k_{1}+k_{2})^{2},m_{1}^{2},m_{2}^{2},m_{3}^{2}\right)$ ), and the two-point function $B_{0}\equiv B_{0}\left(k^{2},m_{4}^{2},m_{3}^{2}\right)$. Since $B_{0}\left(k^{2},m_{4}^{2},m_{3}^{2}\right)$ does not depend on either dispersive or Feynman parameters, we can evaluate dispersive integral multiplied by $B_{0}\left(k^{2},m_{4}^{2},m_{3}^{2}\right)$ analytically. As a result, the first term in Eq. (71) cancels out with dispersive integration times $B_{0}\left(k^{2},m_{4}^{2},m_{3}^{2}\right)$. Final two-loop result has a rather simple form:
$U^{(2)}=-\frac{1}{2\pi}\int\limits_{0}^{1}{\rm d}x\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}\frac{{\rm d}s}{s}\left(s+m_{5}^{2}-m_{12x}^{2}\right)\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}C_{0}.$ (72)
Three-point function can also be written analytically:
$C_{0}=$ $\frac{1}{s-m_{43x}^{2}}\left[x\mathfrak{\ Disc}\left(x^{2}k^{2},m_{3}^{2},s\right)+\bar{x}\ \mathfrak{Disc}\left(\bar{x}^{2}k^{2},m_{3}^{2},s\right)-\mathfrak{Disc}\left(k^{2},m_{3}^{2},m_{4}^{2}\right)\right]$
$+$ $\frac{1}{2x\bar{x}k^{2}}\left(\ln\frac{m_{3}^{2}}{s}-x\ln\frac{m_{3}^{2}}{m_{4}^{2}}\right).$ (73)
Here, $m_{43x}^{2}=xm_{4}^{2}+\bar{x}m_{3}^{2}-x\bar{x}k^{2}$ and $\mathfrak{Disc}\left(k^{2},m_{1}^{2},m_{2}^{2}\right)$ is a discontinuity of the two-point function, which contains branch cut from $(m_{1}+m_{2})^{2}$ to infinity and has the following structure:
$\mathfrak{Disc}\left(k^{2},m_{1}^{2},m_{2}^{2}\right)$ $=\frac{\sqrt{-\Delta\left(k^{2},m_{1}^{2},m_{2}^{2}\right)}}{k^{2}}\ln\left(\frac{m_{1}^{2}+m_{2}^{2}-k^{2}+\sqrt{-\Delta\left(k^{2},m_{1}^{2},m_{2}^{2}\right)}}{2m_{1}m_{2}}\right).$ (74)
At this point, using Eqs. (72)–(74), we can reproduce numerical results for the two-loop graph $U^{(2)}$ in Fig. 6. To make a comparison to the earlier works (Bohm, ) and (AA1, ) we will use $m_{1}=2.0$ GeV, $m_{2}=1.0$ GeV, $m_{3}=4.0$ GeV, $m_{4}=5.0$ GeV and $m_{5}=3.0$ GeV.
$k^{2}$ (GeV2) This work $\Delta t$ (sec) (AA1, ) $\Delta t$ (sec) (Bohm, ) (Table 1) $\Delta t$ (sec)
50.0-50.0 0.08295-0.08295 $1.0$ 0.08296-0.08296 $75.0$ - -
10.0-10.0 0.18399-0.18399 $0.7$ 0.18399-0.18399 $22.0$ - -
5.0-5.0 0.22180-0.22180 $0.7$ 0.22178-0.22178 $17.0$ - -
1.0-1.0 0.26923-0.26923 $0.7$ 0.26919-0.26919 $8.0$ - -
0.5-0.5 0.27704-0.27704 $0.7$ 0.27712-0.27712 $9.0$ - -
0.1-0.1 0.28372-0.28372 $0.7$ 0.28360-0.28360 $9.0$ - -
$0.1$ 0.28723-0.28723 $0.7$ 0.28714-0.28714 $9.0$ 0.28724-0.28724 $0.6$
$0.5$ 0.29458-0.29458 $0.7$ 0.29443-0.29443 $9.0$ 0.29459-0.29459 $0.7$
$1.0$ 0.30451-0.30451 $0.7$ 0.30449-0.30449 $10.0$ 0.30452-0.30452 $0.7$
$5.0$ 0.45250-0.45250 $0.8$ 0.45230-0.45230 $14.0$ 0.45252-0.45252 $0.7$
$10.0$ $8807-0.35309{\rm i} 4.04.0$ $0.48810-0.35318{\rm i}$ $30.0$ $0.48815-0.35322{\rm i}$ $0.7$
$50.0$ $0.17390-0.11804{\rm i}$ $76.0$ $0.17335-0.11781{\rm i}$ $1120.0$ $0.17390-0.11808{\rm i}$ $1.4$
Table 1: Numerical results for $U^{(2)}$ (Fig. 6) calculated from Eq. (72) and compared to (AA1, ) and (Bohm, ). The masses are $m_{1}=2.0$ GeV, $m_{2}=1.0$ GeV, $m_{3}=4.0$ GeV, $m_{4}=5.0$ GeV and $m_{5}=3.0$ GeV.
For the numerical integration we shifted $k^{2}$ and all the masses by ${\rm i}\cdot 10^{-16}$ to remove singular behavior at the poles in Eq. (72). Numerical results in Table 1 are in a very good agreement if compared to previously obtained values in (AA1, ) and (Bohm, ). It is worth noting that in this work, as well as in Ref (AA1, ), Mathematica was used to complete numerical integration using GlobalAdaptive method. In Ref. (Bohm, ), QUADPACK routine was applied for numerical integration.
V Conclusion
In this paper, we continue the development of the dispersive approach for the calculation of multi-loop Feynman diagrams. This study builds upon our previous work, where we introduced a general framework based on representing multi-point Passarino-Veltman functions in a two-point function basis, thereby allowing the replacement of sub-loop diagrams by effective propagators. In the present work, we extend this framework by employing shifted space-time dimensions in the tensor decomposition of the sub-loops, together with recurrence relations that systematically lower both the dimensionality and the powers of propagators. These relations algebraically minimize the number of basic dispersive integrals required for numerical evaluation. Furthermore, the complexity of the resulting expressions can be reduced by subtracting a finite number of terms from the small-momentum expansion, which significantly improves the convergence of the dispersion integrals. Compared to the differentiation-based approach with respect to internal masses used in our earlier study, this algebraic reduction scheme proves substantially more efficient numerically, reducing computation time and enabling the treatment of more complex topologies. Our method complements recent advances in two-loop electroweak calculations by providing a semi-analytical pathway that combines the dispersive representation of sub-loops, dimension-recurrence identities, and the two-point-basis decomposition. Since obtaining fully analytic results for general two-loop, multi-leg electroweak amplitudes remains exceptionally challenging, our approach offers a scalable and robust alternative. Instead of pursuing closed-form solutions for each diagram, we transform the problem into a compact set of well-behaved dispersive integrals amenable to stable numerical evaluation, enabling precision predictions directly applicable to current and upcoming experiments such as MOLLER, P2, and Belle II. Looking ahead, this framework establishes a solid foundation for the automation of multi-loop calculations in a dispersive representation. The next steps will involve implementing the reduction and integration algorithms into a numerical library optimized for precision electroweak observables and extending the method to full two-loop amplitudes with multiple mass scales. Such developments will enable comprehensive, ab-initio predictions for a broad class of processes relevant to the ongoing and future precision programs at JLab, MESA, and KEK, bridging the gap between analytical theory and phenomenological applications. Ultimately, the goal is not merely the refinement of individual calculations but the construction of a predictive, ab-initio framework capable of interpreting deviations in upcoming experiments as definitive signals of new physics. In this broader context, the ongoing development of the dispersive approach, along with canonical two-loop methods and numerical integration tools, represents an essential component of the global precision-physics enterprise.
Acknowledgements.
This work was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC).
Appendix A: Recurrence relations
According to the general notation (1) and (3) we define the scalar two-point integrals as
$J^{(2)}(D;\nu_{1},\nu_{2})=\int\frac{{\rm d}^{D}q}{\left[q^{2}-m_{1}^{2}\right]^{\nu_{1}}\left[(k+q)^{2}-m_{2}^{2}\right]^{\nu_{2}}}\;.$ (75)
Using the integration by parts technique Tkachov ; ChT we get the following recurrence relations for these integrals BDS :
${\mathbf{1}}^{+}J^{(2)}(D;\nu_{1},\nu_{2})$ $\!\!\!\!=\!\!\!\!$ $\frac{1}{\nu_{1}\Delta}\left\{\left[(k^{2}-m_{1}^{2})(D-\nu_{1}-2\nu_{2})+m_{2}^{2}(D-3\nu_{1})\right]\right$. (76)
$\hskip 28.45274pt\left.-2\nu_{2}m_{2}^{2}{\mathbf{1}}^{-}{\mathbf{2}}^{+}-\nu_{1}(k^{2}-m_{1}^{2}-m_{2}^{2}){\mathbf{1}}^{+}{\mathbf{2}}^{-}\right\}J^{(2)}(D;\nu_{1},\nu_{2}),$
${\mathbf{2}}^{+}J^{(2)}(D;\nu_{1},\nu_{2})$ $\!\!\!\!=\!\!\!\!$ $\frac{1}{\nu_{2}\Delta}\left\{\left[(k^{2}-m_{2}^{2})(D-2\nu_{1}-\nu_{2})+m_{1}^{2}(D-3\nu_{2})\right]\right$. (77)
$\hskip 28.45274pt\left.-2\nu_{1}m_{1}^{2}{\mathbf{1}}^{+}{\mathbf{2}}^{-}-\nu_{2}(k^{2}-m_{1}^{2}-m_{2}^{2}){\mathbf{1}}^{-}{\mathbf{2}}^{+}\right\}J^{(2)}(D;\nu_{1},\nu_{2}),\hskip 19.91692pt$
with ${\mathbf{1}}^{\pm}J^{(2)}(D;\nu_{1},\nu_{2})=J^{(2)}(D;\nu_{1}\pm 1,\nu_{2})$, ${\mathbf{2}}^{\pm}J^{(2)}(D;\nu_{1},\nu_{2})=J^{(2)}(D;\nu_{1},\nu_{2}\pm 1)$, and
$\Delta\equiv\Delta(m_{1}^{2},m_{2}^{2},k^{2})$ $=$ $2k^{2}m_{1}^{2}+2k^{2}m_{2}^{2}+2m_{1}^{2}m_{2}^{2}-(k^{2})^{2}-m_{1}^{4}-m_{2}^{4}$ (78)
$=$ $4m_{1}^{2}m_{2}^{2}-(k^{2}-m_{1}^{2}-m_{2}^{2})^{2}$
$=$ $-\left[k^{2}-(m_{1}+m_{2})^{2}\right]\;\left[k^{2}-(m_{1}-m_{2})^{2}\right]$
$=$ $-\lambda(m_{1}^{2},m_{2}^{2},k^{2})\;,$
where $\lambda(m_{1}^{2},m_{2}^{2},k^{2})$ is the standard notation for the Källen function. Note that the sum of the indices $\nu_{1}$ and $\nu_{2}$ on the r.h.s. of Eqs. (76)–(77) is less by one than their sum on the l.h.s. Therefore, by using these relations all the integrals with higher integer $\nu$’s can be expressed in terms of the integral $J^{(2)}(D;1,1)$ and the massive tadpoles
$J^{(2)}(D;\nu_{1},0)$ $\!\!=\!\!$ ${\rm i}^{1-2\nu_{1}}\pi^{D/2}\frac{\Gamma(\nu_{1}\!-\!D/2)}{\Gamma(\nu_{1})}(m_{1}^{2})^{D/2-\nu_{1}}=(-m_{1}^{2})^{1-\nu_{1}}\frac{\Gamma(\nu_{1}-D/2)}{\Gamma(\nu_{1})\Gamma(1\!-\!D/2)}J^{(2)}(D;1,0),$ (79)
$J^{(2)}(D;0,\nu_{2})$ $\!\!=\!\!$ ${\rm i}^{1-2\nu_{2}}\pi^{D/2}\frac{\Gamma(\nu_{2}\!-\!D/2)}{\Gamma(\nu_{2})}(m_{2}^{2})^{D/2-\nu_{2}}=(-m_{2}^{2})^{1-\nu_{2}}\frac{\Gamma(\nu_{2}-D/2)}{\Gamma(\nu_{2})\Gamma(1\!-\!D/2)}J^{(2)}(D;0,1).$ (80)
Furthermore, to bring the shifted values of the space-time dimension $D$ back to $4-2\varepsilon$ (or $2-2\varepsilon$) we can use the following relation, which can be obtained by using the geometrical approach DD-JMP ; D-NIM or the functional relations Tarasov97 :
$J^{(2)}(D+2;\nu_{1},\nu_{2})$ $\!\!\!=\!\!\!$ $-\frac{\pi}{2k^{2}(D-\nu_{1}-\nu_{2}+1)}\left[\Delta J^{(2)}(D;\nu_{1},\nu_{2})\right$. (81)
$\hskip 113.81102pt+(k^{2}+m_{1}^{2}-m_{2}^{2})J^{(2)}(D;\nu_{1},\nu_{2}-1)$
$\hskip 113.81102pt\left.+(k^{2}-m_{1}^{2}+m_{2}^{2})J^{(2)}(D;\nu_{1}-1,\nu_{2})\right]\;.\hskip 14.22636pt$
In particular, for $\nu_{1}=\nu_{2}=1$ Eq. (81) yields
$J^{(2)}(D+2;1,1)$ $\!\!\!=\!\!\!$ $-\frac{\pi}{2k^{2}(D-1)}\left[\Delta J^{(2)}(D;1,1)\right$. (82)
$\left.+(k^{2}+m_{1}^{2}-m_{2}^{2})J^{(2)}(D;1,0)+(k^{2}-m_{1}^{2}+m_{2}^{2})J^{(2)}(D;0,1)\right]\;.\hskip 14.22636pt$
To deal with the occurring tadpole integrals we can use the following formulae (which follow from Eqs. (79) and (80)):
$J^{(2)}(D+2j;1,0)$ $=$ $\pi^{j}(m_{1}^{2})^{j}\frac{\Gamma(1-D/2-j)}{\Gamma(1-D/2)}J^{(2)}(D;1,0),$ (83)
$J^{(2)}(D+2j;0,1)$ $=$ $\pi^{j}(m_{2}^{2})^{j}\frac{\Gamma(1-D/2-j)}{\Gamma(1-D/2)}J^{(2)}(D;0,1).$ (84)
Using these relations we can express any integral $J^{(2)}(D+2j;\nu_{1},\nu_{2})$ (with non-negative integer $j$, $\nu_{1}$ and $\nu_{2}$) in terms of three integrals, $J^{(2)}(D;1,1)$, $J^{(2)}(D;1,0)$ and $J^{(2)}(D;0,1)$.
Usually the recurrence w.r.t. to the space-time dimension $D$ stops at $D=4-2\varepsilon$. However, as an option, we can also use Eq. (82) one more time, to reduce $J^{(2)}(4-2\varepsilon;1,1)$ to $J^{(2)}(2-2\varepsilon;1,1)$ (which is UV-finite as $\varepsilon\to\penalty 10000\ 0$):
$J^{(2)}(4-2\varepsilon;1,1)$ $\!\!\!=\!\!\!$ $-\frac{\pi}{2k^{2}(1-2\varepsilon)}\left[\Delta J^{(2)}(2-2\varepsilon;1,1)\right$. (85)
$\left.+(k^{2}\!+\!m_{1}^{2}\!-\!m_{2}^{2})J^{(2)}(2-2\varepsilon;1,0)+(k^{2}\!-\!m_{1}^{2}\!+\!m_{2}^{2})J^{(2)}(2-2\varepsilon;0,1)\right]\;,\hskip 22.76219pt$
and then use $J^{(2)}(2-2\varepsilon;1,1)$, $J^{(2)}(2-2\varepsilon;1,0)$ and $J^{(2)}(2-2\varepsilon;0,1)$ as the master integrals.
Combining Eqs. (76)–(77) and (81) one can get another pair of useful relations Tarasov97 ,
$J^{(2)}(D+2;\nu_{1}+1,\nu_{2})$ $=$ $-\frac{\pi}{2\nu_{1}k^{2}}\left[(k^{2}-m_{1}^{2}+m_{2}^{2})J^{(2)}(D;\nu_{1},\nu_{2})\right$. (86)
$\hskip 45.5244pt\left.+J^{(2)}(D;\nu_{1},\nu_{2}-1)-J^{(2)}(D;\nu_{1}-1,\nu_{2})\right],$
$J^{(2)}(D+2;\nu_{1},\nu_{2}+1)$ $=$ $-\frac{\pi}{2\nu_{2}k^{2}}\left[(k^{2}+m_{1}^{2}-m_{2}^{2})J^{(2)}(D;\nu_{1},\nu_{2})\right$. (87)
$\hskip 45.5244pt\left.-J^{(2)}(D;\nu_{1},\nu_{2}-1)+J^{(2)}(D;\nu_{1}-1,\nu_{2})\right].$
They can be used to simultaneously reduce one of the indices ($\nu_{1}$ or $\nu_{2}$) and the space-time dimension $D$. A nice property of Eqs. (86)–(87) is the absence of $\Delta$ in the denominators.
Appendix B: $\varepsilon$-expansion of the master integral
In general, the expansion of the master integral $J^{(2)}(4-2\varepsilon;1,1)$ is known to an arbitrary order in $\varepsilon$ Crete ; D-ep ; DK1 . Keeping terms up to the order $\varepsilon$ we get
$J^{(2)}\left(4-2\varepsilon;1,1\right)$ $\!\!=\!\!$ $\frac{\mbox{i}\pi^{2-\varepsilon}\Gamma(1+\varepsilon)}{2(1-2\varepsilon)}\Biggl\{\frac{(m_{1}^{2})^{-\varepsilon}+(m_{2}^{2})^{-\varepsilon}}{\varepsilon}+\frac{m_{1}^{2}-m_{2}^{2}}{\varepsilon k^{2}}\left[(m_{1}^{2})^{-\varepsilon}-(m_{2}^{2})^{-\varepsilon}\right]$ (88)
$\hskip 0.0pt+\left[1+\varepsilon\ln\left(\frac{k^{2}}{\Delta}\right)\right]F_{1}+2\varepsilon F_{2}+{\cal{O}}(\varepsilon^{2})\Biggr\}\;,\hskip 22.76219pt$
where $\Delta\equiv\Delta(m_{1}^{2},m_{2}^{2},k^{2})$ is defined in Eq. (78). For the integral in $2-2\varepsilon$ dimensions we get (e.g., using Eq. (85))
$J^{(2)}\left(2-2\varepsilon;1,1\right)=-\mbox{i}\pi^{1-\varepsilon}\Gamma(1+\varepsilon)\;\frac{k^{2}}{\Delta}\left\{\left[1+\varepsilon\ln\left(\frac{k^{2}}{\Delta}\right)\right]F_{1}+2\varepsilon F_{2}+{\cal{O}}(\varepsilon^{2})\right\}\;.\hskip 22.76219pt$ (89)
Between the pseudothreshold and the threshold, when $(m_{1}-m_{2})^{2}\leq k^{2}\leq(m_{1}+m_{2})^{2}$ and $\Delta\geq 0$, the functions $F_{i}$ can be presented as
$F_{i}=\frac{\sqrt{\Delta}}{k^{2}}\sum\limits_{i=1}^{2}\left[{\mbox{Ls}}_{i}\left(\pi\right)-{\mbox{Ls}}_{i}\left(2\tau^{\prime}_{0i}\right)\right]\;,$ (90)
where
$\cos\tau^{\prime}_{01}=\frac{m_{1}^{2}-m_{2}^{2}+k^{2}}{2m_{1}\sqrt{k^{2}}},\qquad\cos\tau^{\prime}_{02}=\frac{m_{2}^{2}-m_{1}^{2}+k^{2}}{2m_{2}\sqrt{k^{2}}},$ (91)
and the log-sine integrals are defined as
${\mbox{Ls}}_{j}\left(\theta\right)\equiv-\int_{0}^{\theta}\mbox{d}\theta^{\prime}\;\ln^{j-1}\left|2\sin\frac{\theta}{2}\right|\;.$ (92)
In particular, ${\mbox{Ls}}_{1}\left(\theta\right)=-\theta$, and ${\mbox{Ls}}_{2}\left(\theta\right)={\mbox{Cl}}_{2}\left(\theta\right)$, where
${\mbox{Cl}}_{2}\left(\theta\right)=\frac{1}{2\mbox{i}}\left[{\mbox{Li}}_{2}\left(e^{\rm{i}\theta}\right)-{\mbox{Li}}_{2}\left(e^{-\rm{i}\theta}\right)\right]$ (93)
is the Clausen function. Therefore,
$F_{1}$ $=$ $\frac{\sqrt{\Delta}}{k^{2}}\;\sum\limits_{i=1}^{2}\left[{\mbox{Ls}}_{1}\left(\pi\right)-{\mbox{Ls}}_{1}\left(2\tau^{\prime}_{0i}\right)\right]=-2\frac{\sqrt{\Delta}}{k^{2}}\;\arccos{\left(\frac{m_{1}^{2}+m_{2}^{2}-k^{2}}{2m_{1}m_{2}}\right)}\;,$ (94)
$F_{2}$ $=$ $\frac{\sqrt{\Delta}}{k^{2}}\sum\limits_{i=1}^{2}\left[{\mbox{Ls}}_{2}\left(\pi\right)-{\mbox{Ls}}_{2}\left(2\tau^{\prime}_{0i}\right)\right]=-\frac{\sqrt{\Delta}}{k^{2}}\left[{\mbox{Cl}}_{2}\left(2\tau^{\prime}_{01}\right)+{\mbox{Cl}}_{2}\left(2\tau^{\prime}_{02}\right)\right]\;.$ (95)
In other regions (where $\Delta<0$) one can use analytic continuation. This process was described in Ref. DK1 , at any order in $\varepsilon$. Introducing variables $z_{i}=e^{{\rm i}\sigma\theta_{i}}$, such that $\theta_{i}=2\tau^{\prime}_{0i}$, $\sigma=\pm 1$, we get
${\rm i}\sigma\left[{\mbox{Ls}}_{1}\left(\pi\right)-{\mbox{Ls}}_{1}\left(\theta_{i}\right)\right]$ $=$ $\ln(-z_{i})\;,$ (96)
${\rm i}\sigma\left[{\mbox{Ls}}_{2}\left(\pi\right)-{\mbox{Ls}}_{2}\left(\theta_{i}\right)\right]$ $=$ $-\frac{{}_{1}}{{}^{2}}\left[{\mbox{Li}}_{2}\left(z_{i}\right)-{\mbox{Li}}_{2}\left(1/z_{i}\right)\right]\;.$ (97)
In our case, the variables $z_{1}$ and $z_{2}$ can be presented as
$z_{1}=\frac{k^{2}+m_{1}^{2}-m_{2}^{2}+\sqrt{-\Delta}}{k^{2}+m_{1}^{2}-m_{2}^{2}-\sqrt{-\Delta}},\qquad z_{2}=\frac{k^{2}-m_{1}^{2}+m_{2}^{2}+\sqrt{-\Delta}}{k^{2}-m_{1}^{2}+m_{2}^{2}-\sqrt{-\Delta}}.$ (98)
In this way, we get
$F_{1}$ $=$ $-\frac{\sqrt{-\Delta}}{k^{2}}\left[\ln\left(-z_{1}\right)+\ln\left(-z_{2}\right)\right]\;,$ (99)
$F_{2}$ $=$ $-\frac{\sqrt{-\Delta}}{2k^{2}}\left[{\mbox{Li}}_{2}\left(\frac{1}{z_{1}}\right)-{\mbox{Li}}_{2}\left(z_{1}\right)+{\mbox{Li}}_{2}\left(\frac{1}{z_{2}}\right)-{\mbox{Li}}_{2}\left(z_{2}\right)\right]\;.$ (100)
In particular, above the threshold (for $k^{2}>(m_{1}+m_{2})^{2}$, where $z_{1}>1$ and $z_{2}>1$) we can explicitly separate the real and imaginary parts:
$F_{1}$ $=$ $-\frac{\sqrt{-\Delta}}{k^{2}}\left[\ln\left(z_{1}z_{2}\right)+2\rm{i}\pi\right]\;,$ (101)
$F_{2}$ $=$ $-\frac{\sqrt{-\Delta}}{2k^{2}}\left[2{\mbox{Li}}_{2}\left(\frac{1}{z_{1}}\right)+2{\mbox{Li}}_{2}\left(\frac{1}{z_{2}}\right)-\frac{{}_{2}}{{}^{3}}\pi^{2}\!+\!\frac{{}_{1}}{{}^{2}}\ln^{2}{z_{1}}\!+\!\frac{{}_{1}}{{}^{2}}\ln^{2}{z_{2}}+{\rm i}\pi\left(\ln{z_{1}}\!+\!\ln{z_{2}}\right)\right].\hskip 19.91692pt$ (102)
Note that in Eq. (88) we also need to take care of the term
$\ln\left(\frac{k^{2}}{\Delta}\right)F_{1}\Rightarrow\left[\ln\left(\frac{k^{2}}{-\Delta}\right)-{\rm i}\pi\right]F_{1}\;.$ (103)
In this way, we get the following result for the imaginary part of ${\rm i}^{-1}J^{(2)}\left(4-2\varepsilon;1,1\right)$ above the threshold:
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(4-2\varepsilon;1,1)\right]$ $=$ $-\frac{\pi^{2-\varepsilon}\Gamma(1+\varepsilon)}{(1-2\varepsilon)}\frac{\pi\sqrt{-\Delta}}{k^{2}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}$ (104)
$=$ $-\pi^{2-\varepsilon}e^{-\gamma\varepsilon}\;\frac{\pi\sqrt{-\Delta}}{k^{2}}\Biggl\{1+2\varepsilon+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}\;.\hskip 28.45274pt$
For $J^{(2)}(2-2\varepsilon;1,1)$ we get (e.g., using Eq. (85)):
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]$ $=$ $-2\pi^{1-\varepsilon}\Gamma(1+\varepsilon)\;\frac{\pi}{\sqrt{-\Delta}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}$ (105)
$=$ $-2\pi^{1-\varepsilon}e^{-\gamma\varepsilon}\;\frac{\pi}{\sqrt{-\Delta}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}\;.\hskip 28.45274pt$
One can also obtain the result for an arbitrary $\varepsilon$ (see in DR ; DD-JPA )
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\frac{\pi}{\sqrt{-\Delta}}\left(\frac{k^{2}}{-\Delta}\right)^{\varepsilon}\;.$ (106)
Appendix C: Small-$k^{2}$ expansion of the two-point function
For small $k^{2}$ and arbitrary $m_{1}$ and $m_{2}$, using Eq. (20) of Ref. BD-TMF we get (for unit powers of propagators)
$J^{(2)}(D;1,1)$ $\!\!=\!\!$ $-{\rm i}\pi^{D/2}(m_{2}^{2})^{D/2-2}\Gamma(1-D/2)$ (107)
$\times\Biggl\{F_{4}\left(1,2-D/2;D/2,2-D/2\Bigl|\frac{k^{2}}{m_{2}^{2}},\frac{m_{1}^{2}}{m_{2}^{2}}\right)$
$\hskip 8.53581pt-\left(\frac{m_{1}^{2}}{m_{2}^{2}}\right)^{D/2-1}F_{4}\left(1,2-D/2;D/2,2-D/2\Bigl|\frac{k^{2}}{m_{2}^{2}},\frac{m_{1}^{2}}{m_{2}^{2}}\right)\;,\hskip 22.76219pt$
where
$F_{4}(a,b;c,d|x,y)=\sum\limits_{j_{1}=0}^{\infty}\sum\limits_{j_{2}=0}^{\infty}\frac{(a)_{j_{1}+j_{2}}(b)_{j_{1}+j_{2}}}{(c)_{j_{1}}(d)_{j_{2}}}\frac{x^{j_{1}}y^{j_{2}}}{j_{1}!j_{2}!}$ (108)
is Appell’s hypergeometric function of two variables. The sum over $j_{2}$ produces Gauss hypergeometric function,
$F_{4}(a,b;c,d|x,y)=\sum\limits_{j=0}^{\infty}\frac{x^{j}}{j!}\frac{(a)_{j}(b)_{j}}{(c)_{j}}\;{}_{2}F_{1}\left(\left.\begin{array}[]{c}a+j,b+j\\ d\end{array}\right|y\right)\;.$ (109)
Therefore, we can express the integral (107) as
$J^{(2)}(D;1,1)$ $\!\!=\!\!$ $-\frac{{\rm i}\pi^{D/2}}{m_{2}^{2}}\sum\limits_{j=0}^{\infty}\left(\frac{k^{2}}{m_{2}^{2}}\right)^{j}\frac{1}{(D/2)_{j}}$ (115)
$\times\Biggl\{(m_{2}^{2})^{D/2-1}(2-D/2)_{j}\;{}_{2}F_{1}\left(\left.\begin{array}[]{c}1+j,2-D/2+j\\ 2-D/2\end{array}\right|\frac{m_{1}^{2}}{m_{2}^{2}}\right)$
$\hskip 14.22636pt-(m_{1}^{2})^{D/2-1}(D/2)_{j}\;{}_{2}F_{1}\left(\left.\begin{array}[]{c}1+j,D/2+j\\ D/2\end{array}\right|\frac{m_{1}^{2}}{m_{2}^{2}}\right)\Biggr\}\;.$
The occurring $_{2}F_{1}$ functions can be transformed into truncating $_{2}F_{1}$ functions (see, e.g., Eq. 7.3.1.26 of Ref. PBM3 ),
${}_{2}F_{1}\left(\left.\begin{array}[]{c}1+j,1+\alpha+j\\ 1+\alpha\end{array}\right|z\right)$ $=$ $(1-z)^{-1-2j}{}_{2}F_{1}\left(\left.\begin{array}[]{c}-j,\alpha-j\\ 1+\alpha\end{array}\right|z\right)\;,$ (120)
where $\alpha=1-D/2$ in the first case and $\alpha=D/2-1$ in the second case. The resulting finite sums can be written as
$_{2}F_{1}\left(\left.\begin{array}[]{c}-j,\alpha-j\\ 1+\alpha\end{array}\right|z\right)=\sum\limits_{l=0}^{j}\frac{(\alpha-j)_{l}}{(\alpha+1)_{l}}\;\frac{j!}{l!(j-l)!}(-z)^{l},$ (121)
where the ratio of factorials is nothing but the binomial coefficient.
Using Eq. (121) we get
$J^{(2)}(D;1,1)$ $\!\!=\!\!$ $-{\rm i}\pi^{D/2}\Gamma\left(1-\frac{D}{2}\right)\sum\limits_{j=0}^{\infty}\frac{(k^{2})^{j}}{(D/2)_{j}}\;\frac{1}{(m_{2}^{2}-m_{1}^{2})^{1+2j}}$ (122)
$\times\Biggl\{(m_{2}^{2})^{D/2-1}(2-D/2)_{j}\;\sum\limits_{l=0}^{j}\frac{(1-D/2-j)_{l}}{(2-D/2)_{l}}\;\frac{j!}{l!(j-l)!}(-m_{1}^{2})^{l}(m_{2}^{2})^{j-l}$
$\hskip 14.22636pt-(m_{1}^{2})^{D/2-1}(D/2)_{j}\;\sum\limits_{l=0}^{j}\frac{(D/2-1-j)_{l}}{(D/2)_{l}}\;\frac{j!}{l!(j-l)!}(-m_{1}^{2})^{l}(m_{2}^{2})^{j-l}\Biggr\}\;.\hskip 22.76219pt$
Applying well-known transformations of the Pochhammer symbols, we arrive at an explicitly symmetric result:
$J^{(2)}(D;1,1)$ $\!\!=\!\!$ $-{\rm i}\pi^{D/2}\Gamma\left(1-\frac{D}{2}\right)\sum\limits_{j=0}^{\infty}(k^{2})^{j}\;\frac{(2-D/2)_{j}}{(m_{2}^{2}-m_{1}^{2})^{1+2j}}$ (123)
$\times\Biggl\{(m_{2}^{2})^{D/2-1}\sum\limits_{l=0}^{j}\frac{j!}{l!(j-l)!}\;\frac{(m_{2}^{2})^{l}(m_{1}^{2})^{j-l}}{(D/2)_{l}(2-D/2)_{j-l}}\;$
$\hskip 14.22636pt-(m_{1}^{2})^{D/2-1}\sum\limits_{l=0}^{j}\frac{j!}{l!(j-l)!}\;\frac{(m_{1}^{2})^{l}(m_{2}^{2})^{j-l}}{(D/2)_{l}(2-D/2)_{j-l}}\;\Biggr\}\;.\hskip 22.76219pt$
This result can be also presented in terms of the tadpole integrals $J^{(2)}(D;1,0)$ and $J^{(2)}(D;0,1)$:
$J^{(2)}(D;1,1)$ $\!\!=\!\!$ $\sum\limits_{j=0}^{\infty}(k^{2})^{j}\;\frac{(2-D/2)_{j}}{(m_{2}^{2}-m_{1}^{2})^{1+2j}}$ (124)
$\times\Biggl\{J^{(2)}(D;0,1)\;\sum\limits_{l=0}^{j}\frac{j!}{l!(j-l)!}\;\frac{(m_{2}^{2})^{l}(m_{1}^{2})^{j-l}}{(D/2)_{l}(2-D/2)_{j-l}}\;$
$\hskip 14.22636pt-J^{(2)}(D;1,0)\;\sum\limits_{l=0}^{j}\frac{j!}{l!(j-l)!}\;\frac{(m_{1}^{2})^{l}(m_{2}^{2})^{j-l}}{(D/2)_{l}(2-D/2)_{j-l}}\;\Biggr\}\;.\hskip 22.76219pt$
It is easy to check that the limit $m_{1}=m_{2}\equiv m$ is regular. In this case, Eq. (17) of Ref. BD-TMF yields
$J^{(2)}(D;1,1)\Bigr|_{k^{2}\to 0,\;m_{1}=m_{2}\equiv m}$ $=$ ${\rm i}\pi^{D/2}\;(m^{2})^{D/2-2}\Gamma\left(2-\frac{D}{2}\right)\;{}_{2}F_{1}\left(\left.\begin{array}[]{c}1,2-D/2\\ 3/2\end{array}\right|\frac{k^{2}}{4m^{2}}\right)$ (127)
$=$ ${\rm i}\pi^{D/2}\;(m^{2})^{D/2-2}\Gamma\left(2-\frac{D}{2}\right)\;\sum\limits_{j=0}^{\infty}\frac{(2-D/2)_{j}}{(3/2)_{j}}\;\left(\frac{k^{2}}{4m^{2}}\right)^{j}.\hskip 22.76219pt$ (128)
Using Eq. (122), we have checked that the first 20 terms of its $k^{2}$-expansion in the limit $m_{2}\to\penalty 10000\ m_{1}$ are the same as in Eq. (127).
Let us also consider the case $m_{1}=0$, $m_{2}\equiv m$. In this case, using Eq. (10) of Ref. BD-TMF , we get
$J^{(2)}(D;1,1)\Bigr|_{k^{2}\to 0,\;m_{1}=0,\;m_{2}\equiv m}$ $\!\!=$ $-{\rm i}\pi^{D/2}\;(m^{2})^{D/2-2}\Gamma\left(1-\frac{D}{2}\right)\;{}_{2}F_{1}\left(\left.\begin{array}[]{c}1,2-D/2\\ D/2\end{array}\right|\frac{k^{2}}{m^{2}}\right)$ (131)
$\!\!=$ $-{\rm i}\pi^{D/2}\;(m^{2})^{D/2-2}\Gamma\left(1-\frac{D}{2}\right)\;\sum\limits_{j=0}^{\infty}\frac{(2\!-\!D/2)_{j}}{(D/2)_{j}}\;\left(\frac{k^{2}}{m^{2}}\right)^{j}.\hskip 25.60747pt$ (132)
Let us now consider the limit $m_{1}=0$, $m_{2}\equiv m$ in Eq. (122). The third line (containing $(m_{1}^{2})^{D/2-1}$) should be omitted because it corresponds to a massless tadpole. In the sum on the second line we only need to keep the term with $l=j$ because all the others vanish. As a result we get
${\rm i}\pi^{D/2}\;(m^{2})^{D/2-2}\Gamma\left(1-\frac{D}{2}\right)\;\sum\limits_{j=0}^{\infty}\frac{(D/2-1-j)_{j}}{(D/2)_{j}}\left(-\frac{k^{2}}{m^{2}}\right)^{j}.$
Transforming the Pochhammer symbol as $(D/2-1-j)_{j}=(-1)^{j}(2-D/2)_{j}$ we reproduce the same result as in Eq. (131).
Appendix D: Special cases of the two-point function
In the special case $m_{1}=0$, $m_{2}\equiv m$ we can use Eqs. (2.24)–(2.25) of Ref. DK1 . In particular, in Eq. (2.25) an arbitrary term of the $\varepsilon$-expansion is presented in terms of Nielsen polylogarithms $S_{a,b}(u)$, with $u\equiv k^{2}/m^{2}$. Taking into account that $S_{a,1}(u)={\mbox{Li}}_{a+1}\left(u\right)$ we get
$J^{(2)}(4-2\varepsilon;1,1)\Bigr|_{m_{1}=0,\;m_{2}\equiv m}$ $=$ ${\rm i}\pi^{2-\varepsilon}(m^{2})^{-\varepsilon}\frac{\Gamma(1+\varepsilon)}{1-2\varepsilon}\Biggl\{\frac{1}{\varepsilon}-\frac{1-u}{2u\varepsilon}\left[(1-u)^{-2\varepsilon}-1\right]$ (133)
$-\frac{\varepsilon(1-u)^{-2\varepsilon}}{u}{\mbox{Li}}_{2}\left(u\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}.$
The threshold corresponds to the point $u=1$ ($k^{2}=m^{2}$). To go beyond the threshold we can use
${\mbox{Li}}_{2}\left(u\right)=-{\mbox{Li}}_{2}\left(\frac{1}{u}\right)+\frac{{}_{1}}{{}^{3}}\pi^{2}-\frac{{}_{1}}{{}^{2}}\ln^{2}{u}-{\rm i}\pi\ln{u}$
(the sign of the imaginary part needs to be fixed, depending on the causal prescription). Using
${\mbox{Li}}_{2}\left(u\right)=-{\mbox{Li}}_{2}\left(1-u\right)+\frac{{}_{1}}{{}^{6}}\pi^{2}-\ln{u}\ln(1-u)$
we can get as many terms of the “on-shell" expansion in powers and logarithms of $1-u=(m^{2}-k^{2})/m^{2}$ as we like. In particular, at $k^{2}=m^{2}$ ($u=1$) we get
$J^{(2)}(4-2\varepsilon;1,1)\Bigr|_{m_{1}=0,\;m_{2}\equiv m,\;k^{2}=m^{2}}={\rm i}\pi^{2-\varepsilon}(m^{2})^{-\varepsilon}\frac{\Gamma(\varepsilon)}{1-2\varepsilon}\;.$ (134)
For $m_{1}=m_{2}=0$ we can use the well-known result
$J^{(2)}(4-2\varepsilon;1,1)\Bigr|_{m_{1}=m_{2}=0}={\rm i}\pi^{2-\varepsilon}(-k^{2})^{-\varepsilon}\;\frac{\Gamma^{2}(1-\varepsilon)\;\Gamma(\varepsilon)}{\Gamma(2-2\varepsilon)}\;.$
For $k^{2}=0$ and arbitrary $m_{1}$ and $m_{2}$ we get (see, e.g., in Ref. BD-TMF , after Eq. (21))
$J^{(2)}(4-2\varepsilon;1,1)\Bigr|_{k^{2}=0}=-{\rm i}\pi^{2-\varepsilon}\Gamma(-1+\varepsilon)\;\frac{(m_{1}^{2})^{1-\varepsilon}-(m_{2}^{2})^{1-\varepsilon}}{m_{1}^{2}-m_{2}^{2}}\;.$ (135)
For $m_{1}=m_{2}\equiv m$ Eq. (135) yields
$J^{(2)}(4-2\varepsilon;1,1)\Bigr|_{k^{2}=0,\;m_{1}=m_{2}\equiv m}={\rm i}\pi^{2-\varepsilon}\Gamma(\varepsilon)\;(m^{2})^{-\varepsilon}\;,$
and for $m_{1}=0$, $m_{2}\equiv m$ we get
$J^{(2)}(4-2\varepsilon;1,1)\Bigr|_{k^{2}=0,\;m_{1}=0,\;m_{2}\equiv m}=-{\rm i}\pi^{2-\varepsilon}\Gamma(-1+\varepsilon)\;(m^{2})^{-\varepsilon}\;.$
Considering the case $m_{1}\equiv m$, $m_{2}=\lambda$, $k^{2}=m^{2}$ and using the general hypergeometric representation, after some transformations we get
$J^{(2)}(4\!-\!2\varepsilon;1,1)\Bigr|_{k^{2}=m^{2},\;m_{1}\equiv m,\;m_{2}\equiv\lambda}$ $\!\!\!=\!\!\!$ ${\rm i}\pi^{2-\varepsilon}\Biggl\{\frac{\Gamma(\varepsilon)(m^{2})^{-\varepsilon}}{1-2\varepsilon}\left(1\!-\!\frac{\lambda^{2}}{2m^{2}}\right){}_{2}F_{1}\Biggl(\begin{array}[]{c}1,\;\varepsilon\\ 1/2\!+\!\varepsilon\end{array}\Biggr|\frac{\lambda^{2}(4m^{2}\!-\!\lambda^{2})}{4m^{4}}\Biggr)$ (141)
$+\Gamma\left(\frac{{}_{3}}{{}^{2}}\right)\Gamma\left(-\frac{{}_{1}}{{}^{2}}+\varepsilon\right)(m^{2})^{-1/2}(\lambda^{2})^{1/2-\varepsilon}\left(1-\frac{\lambda^{2}}{4m^{2}}\right)^{1/2-\varepsilon}$
$+\frac{\Gamma(\varepsilon)}{2m^{2}}(\lambda^{2})^{1-\varepsilon}{}_{2}F_{1}\Biggl(\begin{array}[]{c}1,\;\varepsilon\\ 3/2\end{array}\Biggr|\frac{\lambda^{2}}{4m^{2}}\Biggr)\Biggr\}\;.$
In the limit $\lambda\to 0$ we reproduce Eq. (134).

### v2

@@
 During the following decade, Refs.
-(Hollik1990, ; Denner1993, ) produced comprehensive reviews and explicit one-loop calculations relevant to LEP precision physics, codifying gauge-invariant renormalization schemes and parameter definitions.
+(Hollik1990, ; Denner1993, ) produced comprehensive reviews and explicit one-loop calculations relevant to CERN’s Large Electron-Positron (LEP) collider precision physics, codifying gauge-invariant renormalization schemes and parameter definitions.
 These works became standard references for both theoretical and computational approaches to radiative corrections.
@@
 This formalism remains foundational for modern two-loop reduction algorithms and underpins most symbolic-manipulation packages used today.
-Concurrently, in (PLB91, ) a complementary algebraic framework was developed for reducing tensor Feynman integrals to scalar ones using dimension shifting recurrence relations.
+Concurrently, in (PLB91, ) a complementary algebraic framework was developed for reducing tensor Feynman integrals to scalar ones using dimension shifting recurrence relations (see also (GG, ; BGHPS, ; EGZ, ) and references therein).
 Analytic evaluation of increasingly complex diagrams soon became infeasible due to the proliferation of mass scales, external invariants, and threshold singularities.
-This challenge prompted the development of semi-analytical and numerical approaches such as sector-decomposition methods, differential-equation (DE) systems for master integrals, and dispersion-relation techniques (Lee, ; SmrBook, ; Kotikov, ).
+This challenge prompted the development of semi-analytical and numerical approaches such as sector-decomposition methods, differential-equation (DE) systems for master integrals, and dispersion-relation techniques (Lee, ; SmrBook, ; Kotikov, ; BSch, ; BSSch, ).
 A major step was the differential-equation approach for master integrals, in which systems of linear equations in kinematic invariants are solved either analytically in canonical ($\varepsilon$-form) basis or numerically using series expansions.
 In (Wasser, ; Henn, ) this framework was refined by introducing canonical-basis and uniform-weight formulations, improving both analytic transparency and numerical stability.
-In (AA-1, ; AA-2, ; AA-3, ; AA-4, ; AA-5, ), it was developed a comprehensive two-loop framework, specifically for polarized Møller scattering: an essential channel for future parity-violation experiments.
+A comprehensive two-loop framework was developed in (AA-1, ; AA-2, ; AA-3, ; AA-4, ; AA-5, ), specifically for polarized Møller scattering: an essential channel for future parity-violation experiments.
 The group systematically advanced from reducible two-loop and quadratic one-loop contributions to two-loop irreducible self-energies, vertex and box calculations.
 These results quantified higher-order electroweak effects in parity-violating asymmetries, establishing reliable theoretical uncertainties at the sub-percent level-critical for the forthcoming MOLLER experiment.
-Beyond differential-equation and sector-decomposition strategies, a distinct line of development has been the dispersive approach developed in (AA1, )-(AB19a, ).
+Beyond differential-equation and sector-decomposition strategies, a distinct line of development has been the dispersive approach introduced in (AA1, )-(AB19a, ).
 This methodology expresses multi-point Passarino-Veltman functions in a two-point-function basis, replacing sub-loop insertions with effective propagators represented by dispersion integrals.
@@
 Furthermore, real-experiment implementation often involves acceptance and energy-threshold cuts that significantly affect radiative corrections; the dispersive formulation allows these to be incorporated naturally at the numerical-integration stage.
+An example of an application of the dispersive approach to multi-loop “banana" integrals can be found in Ref.
+ChYCh .
+Another example is the precision studies of the Higgs boson at future colliders, which requires the inclusion of NNLO electroweak corrections.
+In (FR1a, ), the planar and non-planar double-box topologies with multiple massive propagators in the loops are computed numerically by transforming one of the sub-loops using Feynman parametrization and a dispersion representation.
+This was followed by (FR1b, ), which provides a complete calculation of the NNLO electroweak corrections involving closed fermion loops.
 Parallel developments in the phenomenological sector culminated in partial two-loop electroweak predictions for key observables.
 In Refs.
-(Dubovyk1, ) and (Freitas1a, ) produced the full set of fermionic and bosonic two-loop corrections to Z-boson observables, while (Erler, ) addressed hadronic effects in Møller scattering at NNLO.
+(Dubovyk1, ) and (Freitas1a, ) it was produced the full set of fermionic and bosonic two-loop corrections to Z-boson observables, while (Erler, ) addressed hadronic effects in Møller scattering at NNLO.
 More recently, (Schw, ) achieved the analytic evaluation of electroweak double-box integrals relevant for Møller processes.
 These advances collectively establish a robust infrastructure for high-precision electroweak phenomenology.
-A comprehensive review on updated measurements and higher-order theoretical corrections is available at (Erler1, ) and (PDG2024, ).
+The comprehensive review on updated measurements and higher-order theoretical corrections are available at (Erler1, ) and (PDG2024, ).
 Despite impressive progress, complete NNLO electroweak results remain available only for selected processes due to the technical difficulty of two-loop calculations.
@@
 where (see Eq. (109))
-${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]_{s}=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\frac{\pi}{\sqrt{-\Delta_{s}}}\left(\frac{s}{-\Delta_{s}}\right)^{\varepsilon}\;$ (42)
+${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]_{s}=2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\frac{\pi}{\sqrt{-\Delta_{s}}}\left(\frac{s}{-\Delta_{s}}\right)^{\varepsilon}\;$ (42)
 (the subscript $s$ means that we substitute $k^{2}\to s$).
@@
 To proceed, let us use the analytic result given in Eq. (42) to calculate the derivatives
-$\frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}\right\}=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\pi\;\left(\frac{\partial\Delta_{s}}{\partial s}\right)\left(\frac{\partial}{\partial\Delta_{s}}(-\Delta_{s})^{-1/2-\varepsilon}\right)\;,$
-$\frac{\partial}{\partial m_{2}^{2}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\pi\;s^{\varepsilon}\left(\frac{\partial\Delta_{s}}{\partial m_{2}^{2}}\right)\left(\frac{\partial}{\partial\Delta_{s}}(-\Delta_{s})^{-1/2-\varepsilon}\right)\;.$
+$\frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}\right\}=2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\pi\;\left(\frac{\partial\Delta_{s}}{\partial s}\right)\left(\frac{\partial}{\partial\Delta_{s}}(-\Delta_{s})^{-1/2-\varepsilon}\right)\;,$
+$\frac{\partial}{\partial m_{2}^{2}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}=2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\pi\;s^{\varepsilon}\left(\frac{\partial\Delta_{s}}{\partial m_{2}^{2}}\right)\left(\frac{\partial}{\partial\Delta_{s}}(-\Delta_{s})^{-1/2-\varepsilon}\right)\;.$
 Combining these equations we get
@@
 4.
-Figure 4: Numerical results for the four-point function $D_{0}$ ($k_{2}^{2}=-1.5\penalty 10000\ {\rm GeV}^{2}$, $k_{3}^{2}=-2.5\penalty 10000\ {\rm GeV}^{2}$, $k_{4}^{2}=m_{4}^{2}\penalty 10000\ {\rm GeV}^{2}$, $(k_{1}\cdot k_{3})=4.0\penalty 10000\ {\rm GeV}^{2}$, $(k_{2}\cdot k_{3})=-1.0\penalty 10000\ {\rm GeV}^{2}$ $m_{1}=1.5\penalty 10000\ {\rm GeV}$, $m_{2}=0.5\penalty 10000\ {\rm GeV}$, $m_{3}=2.0\penalty 10000\ {\rm GeV}$, $m_{4}=2.5\penalty 10000\ {\rm GeV}$).
+Figure 4: Numerical results for the four-point function $D_{0}$ ($k_{2}^{2}=-1.5\penalty 10000\ {\rm GeV}^{2}$, $k_{3}^{2}=-2.5\penalty 10000\ {\rm GeV}^{2}$, $k_{4}^{2}=m_{4}^{2}$, $(k_{1}\cdot k_{3})=4.0\penalty 10000\ {\rm GeV}^{2}$, $(k_{2}\cdot k_{3})=-1.0\penalty 10000\ {\rm GeV}^{2}$ $m_{1}=1.5\penalty 10000\ {\rm GeV}$, $m_{2}=0.5\penalty 10000\ {\rm GeV}$, $m_{3}=2.0\penalty 10000\ {\rm GeV}$, $m_{4}=2.5\penalty 10000\ {\rm GeV}$).
 The functions $D_{2l,n_{1},n_{2},n_{3}}$ are defined in Eq.
@@
 This particular representation is most valuable for applications in two-loop calculations for any possible particle physics models.
-First, if we consider $\left(j+2\right)$-point Feynman graph as an insertion (index $\left(j+2\right)$ means that we have $j$ number of external and two internal legs in the insertion) into the two-loop topology (see Fig.
+It is worth mentioning that in this chapter we will consider only planar topology for the two-loop graphs.
+In the case of non-planar graphs, we can re-arrange momenta in the loop integral and reduce result to the dispersive representation considered in (AACR, ).
+First, if we consider $\left(j+2\right)$-point Feynman graph as an insertion (index $\left(j+2\right)$ means that we have $j$ number of external and two internal legs in the insertion) into the two-loop planar topology (see Fig.
 5), then polynomial terms in external or second loop momenta will be a part of the numerator algebra and dispersive contribution will be treated as an additional propagator in the second loop integral.
-Figure 5: The two-loop NN-point diagram ($k_{1:j}\equiv k_{1}+\ldots+k_{j}$).
+Figure 5: The two-loop planar NN-point diagram ($k_{1:j}\equiv k_{1}+\ldots+k_{j}$).
 Let us define two-loop integral in the similar way as it was done in Eq. (1):
@@
 $\times\int\int\frac{{\rm d}^{D}q_{1}\;{\rm d}^{D}q_{2}}{\left[q_{1}^{2}-m_{4}^{2}\right]\left[q_{2}^{2}-m_{5}^{2}\right]\left[\left(q_{1}+q_{2}\right)^{2}-m_{1}^{2}\right]\left[\left(q_{1}+q_{2}-k\right)^{2}-m_{2}^{2}\right]\left[\left(q_{1}-k\right)^{2}-m_{3}^{2}\right]},$ (64)
-and first perform integration over the loop momentum $q_{2}$. That results in $J^{(3)}\left(D;1,1,1\right)$:
-$J^{(3)}\left(D;1,1,1\right)=$ $\int\frac{{\rm d}^{D}q_{2}}{\left[q_{2}^{2}-m_{5}^{2}\right]\left[\left(q_{1}+q_{2}\right)^{2}-m_{1}^{2}\right]\left[\left(q_{1}+q_{2}-k\right)^{2}-m_{2}^{2}\right]}.$ (65)
-Using Eq. (II.1), we get
+and first perform integration over the loop momentum $q_{2}$. That results in a three-point integral
+$J^{(3)}\left(D;1,1,1\right)\Bigr|_{\footnotesize{\begin{array}[]{l}p_{1}=0,p_{2}=q_{1},p_{3}=q_{1}-k\\[-8.53581pt] m_{1}^{2}\leftrightarrow m_{5}^{2},m_{2}^{2}\leftrightarrow m_{1}^{2},m_{3}^{2}\leftrightarrow m_{2}^{2}\end{array}}}$ $\!\!\!\!=\int\frac{{\rm d}^{D}q_{2}}{\left[q_{2}^{2}\!-\!m_{5}^{2}\right]\left[\left(q_{1}+q_{2}\right)^{2}\!-\!m_{1}^{2}\right]\left[\left(q_{1}+q_{2}-k\right)^{2}\!-\!m_{2}^{2}\right]}.$ (67)
+Using Eq. (II.1) (with $\bar{x}=1-x$), we get
 $J^{(3)}\left(D;1,1,1\right)=$ $\int\limits_{0}^{1}{\rm d}x\;J^{(2)}\left(D;1,2\right)\Bigr|_{\footnotesize{\begin{array}[]{l}p_{1}=0,p_{2}=q_{1}-\bar{x}k\\[-8.53581pt] m_{1}^{2}\leftrightarrow m_{5}^{2},m_{2}^{2}\leftrightarrow xm_{1}^{2}+\bar{x}m_{2}^{2}-x\bar{x}k^{2}\end{array}}}$ (70)
@@
 which effectively gives us a reduction of the three-point integral to a two-point one.
-Applying recursive approach outlined in Chapter II, we can lower second power of the last propagator in Eq.(69) and arrive to the subtracted $\bar{J}^{(2)}\left(2-2\varepsilon;1,1\right)$ plus terms containing $J^{(2)}\left(2-2\varepsilon;1,0\right)$ and $J^{(2)}\left(2-2\varepsilon;0,1\right)$ (see Eq.
+Here it is assumed that $D=4-2\varepsilon$.
+Applying recursive approach outlined in Chapter II, we can lower second power of the last propagator in Eq.(71) and arrive to the subtracted $\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)$ plus tadpole terms containing $J^{(2)}\left(2-2\varepsilon;1,0\right)$ and $J^{(2)}\left(2-2\varepsilon;0,1\right)$ (see Eq.
 (32)).
-For the subtracted $\bar{J}^{(2)}\left(2-2\varepsilon;1,1\right)$, we can apply dispersive representation stemming from Eqs.
-(41)–(42).
-Finally, for $D=4-2\varepsilon$ and $\varepsilon\to 0$ Eq.
+In this way, Eq.
 (67) can be written in the following form:
-$J^{(3)}\left(4-2\varepsilon;1,1,1\right)$ $=\frac{1}{2}\int\limits_{0}^{1}{\rm d}x\left\{\frac{\rm{i}\pi^{2-\epsilon}}{m_{5}^{2}-m_{12x}^{2}}\ln\frac{m_{12x}^{2}}{m_{5}^{2}}+\left[\left(q_{1}-\bar{x}k\right)^{2}+m_{5}^{2}-m_{12x}^{2}\right]\pi\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)\right\},$ (70)
-where the function $\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)$ is defined in Eq. (41):
-$\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)=\frac{{\rm i}}{\pi}\;\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}}{s\left[\left(q_{1}-\bar{x}k\right)^{2}-s+{\rm i}0\right]}\;.$
-Here, $m_{12x}^{2}=xm_{1}^{2}+\bar{x}m_{2}^{2}-x\bar{x}k^{2}$ and the imaginary part of ${\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)$ at $\varepsilon=0$ has a simple structure:
-$\rm{Im}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}=-\frac{2\pi^{2}}{\sqrt{\left(s-m_{12x}^{2}-m_{5}^{2}\right)^{2}-4m_{12x}^{2}m_{5}^{2}}},$
-In Eq.
-(70) we did not retain linear in $\varepsilon$ terms, since the insertion $J^{(3)}\left(D;1,1,1\right)$, and the entire two-loop integral are UV-finite.
+$J^{(3)}\left(4-2\varepsilon;1,1,1\right)=\frac{1}{2}\int\limits_{0}^{1}{\rm d}x\Biggl\{$ ${\rm i}\pi^{2-\varepsilon}\Gamma(\varepsilon)\;\frac{(m_{5}^{2})^{-\varepsilon}-(m_{12x}^{2})^{-\varepsilon}}{m_{5}^{2}-m_{12x}^{2}}$
+$-\left[\left(q_{1}-\bar{x}k\right)^{2}+m_{5}^{2}-m_{12x}^{2}\right]\pi\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)\Biggr\},$ (72)
+with $m_{12x}^{2}=xm_{1}^{2}+\bar{x}m_{2}^{2}-x\bar{x}k^{2}$.
+For the subtracted $\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)$, we can apply dispersive representation stemming from Eqs.
+(41)–(42),
+$\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)=\frac{{\rm i}}{\pi}\;\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}{\rm d}s\;\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}}{s\left[s-\left(q_{1}-\bar{x}k\right)^{2}-{\rm i}0\right]}\;.$
+Since the insertion $J^{(3)}\left(D;1,1,1\right)$, and the entire two-loop integral are UV-finite, we can consider the limit $\varepsilon\to 0$:
+$J^{(3)}\left(4;1,1,1\right)$ $=\frac{1}{2}\int\limits_{0}^{1}{\rm d}x\left\{\frac{{\rm i}\pi^{2}}{m_{5}^{2}-m_{12x}^{2}}\ln\frac{m_{12x}^{2}}{m_{5}^{2}}-\left[\left(q_{1}-\bar{x}k\right)^{2}+m_{5}^{2}-m_{12x}^{2}\right]\pi\bar{J}_{1}^{(2)}\left(2;1,1\right)\right\}.$ (73)
+Note that at $\varepsilon=0$ the imaginary part of ${\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)$ has a simple structure:
+${\rm Im}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}=\frac{2\pi^{2}}{\sqrt{\left(s-m_{12x}^{2}-m_{5}^{2}\right)^{2}-4m_{12x}^{2}m_{5}^{2}}},$
 At this point, we are ready to complete integration over second loop momentum $q_{1}$:
 $U^{(2)}=$ $\frac{1}{2}\int\limits_{0}^{1}{\rm d}x\Biggl\{\frac{1}{m_{5}^{2}-m_{12x}^{2}}\ln\frac{m_{12x}^{2}}{m_{5}^{2}}\;B_{0}$
-$-\frac{1}{\pi}\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}{\rm d}s\;\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}}{s}\left[\left(s+m_{5}^{2}-m_{12x}^{2}\right)C_{0}+B_{0}\right]\Biggr\},$ (71)
+$+\frac{1}{\pi^{2}}\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}{\rm d}s\;\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}}{s}\left[\left(s+m_{5}^{2}-m_{12x}^{2}\right)C_{0}+B_{0}\right]\Biggr\},$ (74)
 where in the second-loop integration we have used usual PV functions without dispersive representation.
@@
 Final two-loop result has a rather simple form:
-$U^{(2)}=-\frac{1}{2\pi}\int\limits_{0}^{1}{\rm d}x\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}\frac{{\rm d}s}{s}\left(s+m_{5}^{2}-m_{12x}^{2}\right)\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}C_{0}.$ (72)
+$U^{(2)}=\frac{1}{2\pi^{2}}\int\limits_{0}^{1}{\rm d}x\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}\frac{{\rm d}s}{s}\left(s+m_{5}^{2}-m_{12x}^{2}\right)\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}C_{0}.$ (75)
 Three-point function can also be written analytically:
@@
 In particular, above the threshold (for $k^{2}>(m_{1}+m_{2})^{2}$, where $z_{1}>1$ and $z_{2}>1$) we can explicitly separate the real and imaginary parts:
-$F_{1}$ $=$ $-\frac{\sqrt{-\Delta}}{k^{2}}\left[\ln\left(z_{1}z_{2}\right)+2\rm{i}\pi\right]\;,$ (101)
-$F_{2}$ $=$ $-\frac{\sqrt{-\Delta}}{2k^{2}}\left[2{\mbox{Li}}_{2}\left(\frac{1}{z_{1}}\right)+2{\mbox{Li}}_{2}\left(\frac{1}{z_{2}}\right)-\frac{{}_{2}}{{}^{3}}\pi^{2}\!+\!\frac{{}_{1}}{{}^{2}}\ln^{2}{z_{1}}\!+\!\frac{{}_{1}}{{}^{2}}\ln^{2}{z_{2}}+{\rm i}\pi\left(\ln{z_{1}}\!+\!\ln{z_{2}}\right)\right].\hskip 19.91692pt$ (102)
+$F_{1}$ $=$ $-\frac{\sqrt{-\Delta}}{k^{2}}\left[\ln\left(z_{1}z_{2}\right)-2\rm{i}\pi\right]\;,$ (104)
+$F_{2}$ $=$ $-\frac{\sqrt{-\Delta}}{2k^{2}}\left[2{\mbox{Li}}_{2}\left(\frac{1}{z_{1}}\right)+2{\mbox{Li}}_{2}\left(\frac{1}{z_{2}}\right)-\frac{{}_{2}}{{}^{3}}\pi^{2}\!+\!\frac{{}_{1}}{{}^{2}}\ln^{2}{z_{1}}\!+\!\frac{{}_{1}}{{}^{2}}\ln^{2}{z_{2}}-{\rm i}\pi\ln\left(z_{1}z_{2}\right)\right].\hskip 19.91692pt$ (105)
 Note that in Eq. (91) we also need to take care of the term
-$\ln\left(\frac{k^{2}}{\Delta}\right)F_{1}\Rightarrow\left[\ln\left(\frac{k^{2}}{-\Delta}\right)-{\rm i}\pi\right]F_{1}\;.$ (103)
+$\ln\left(\frac{k^{2}}{\Delta}\right)F_{1}\Rightarrow\left[\ln\left(\frac{k^{2}}{-\Delta}\right)+{\rm i}\pi\right]F_{1}\;.$ (106)
 In this way, we get the following result for the imaginary part of ${\rm i}^{-1}J^{(2)}\left(4-2\varepsilon;1,1\right)$ above the threshold:
-${\rm Im}\left[{\rm i}^{-1}J^{(2)}(4-2\varepsilon;1,1)\right]$ $=$ $-\frac{\pi^{2-\varepsilon}\Gamma(1+\varepsilon)}{(1-2\varepsilon)}\frac{\pi\sqrt{-\Delta}}{k^{2}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}$ (104)
-$=$ $-\pi^{2-\varepsilon}e^{-\gamma\varepsilon}\;\frac{\pi\sqrt{-\Delta}}{k^{2}}\Biggl\{1+2\varepsilon+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}\;.\hskip 28.45274pt$
+${\rm Im}\left[{\rm i}^{-1}J^{(2)}(4-2\varepsilon;1,1)\right]$ $=$ $\frac{\pi^{2-\varepsilon}\Gamma(1+\varepsilon)}{(1-2\varepsilon)}\frac{\pi\sqrt{-\Delta}}{k^{2}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}$ (107)
+$=$ $\pi^{2-\varepsilon}e^{-\gamma\varepsilon}\;\frac{\pi\sqrt{-\Delta}}{k^{2}}\Biggl\{1+2\varepsilon+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}\;.\hskip 28.45274pt$
 For $J^{(2)}(2-2\varepsilon;1,1)$ we get (e.g., using Eq. (88)):
-${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]$ $=$ $-2\pi^{1-\varepsilon}\Gamma(1+\varepsilon)\;\frac{\pi}{\sqrt{-\Delta}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}$ (105)
-$=$ $-2\pi^{1-\varepsilon}e^{-\gamma\varepsilon}\;\frac{\pi}{\sqrt{-\Delta}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}\;.\hskip 28.45274pt$
+${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]$ $=$ $2\pi^{1-\varepsilon}\Gamma(1+\varepsilon)\;\frac{\pi}{\sqrt{-\Delta}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}$ (108)
+$=$ $2\pi^{1-\varepsilon}e^{-\gamma\varepsilon}\;\frac{\pi}{\sqrt{-\Delta}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}\;.\hskip 28.45274pt$
 One can also obtain the result for an arbitrary $\varepsilon$ (see in DR ; DD-JPA )
-${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\frac{\pi}{\sqrt{-\Delta}}\left(\frac{k^{2}}{-\Delta}\right)^{\varepsilon}\;.$ (106)
+${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]=2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\frac{\pi}{\sqrt{-\Delta}}\left(\frac{k^{2}}{-\Delta}\right)^{\varepsilon}\;.$ (109)
 Appendix C: Small-$k^{2}$ expansion of the two-point function
@@
 The threshold corresponds to the point $u=1$ ($k^{2}=m^{2}$). To go beyond the threshold we can use
-${\mbox{Li}}_{2}\left(u\right)=-{\mbox{Li}}_{2}\left(\frac{1}{u}\right)+\frac{{}_{1}}{{}^{3}}\pi^{2}-\frac{{}_{1}}{{}^{2}}\ln^{2}{u}-{\rm i}\pi\ln{u}$
-(the sign of the imaginary part needs to be fixed, depending on the causal prescription). Using
+${\mbox{Li}}_{2}\left(u\right)=-{\mbox{Li}}_{2}\left(\frac{1}{u}\right)+\frac{{}_{1}}{{}^{3}}\pi^{2}-\frac{{}_{1}}{{}^{2}}\ln^{2}{u}+{\rm i}\pi\ln{u}.$
+Here the sign of the imaginary part is fixed according to the causal prescription ($k^{2}\leftrightarrow k^{2}+{\rm i}0$). Using
 ${\mbox{Li}}_{2}\left(u\right)=-{\mbox{Li}}_{2}\left(1-u\right)+\frac{{}_{1}}{{}^{6}}\pi^{2}-\ln{u}\ln(1-u)$
@@
 In the limit $\lambda\to 0$ we reproduce Eq. (137).
+Analogous results for special cases of $J^{(2)}(2-2\varepsilon;1,1)$ can be obtained by using Eq. (88).